### v1

# Evolving from Single-modal to Multi-modal Facial Deepfake Detection: A Survey

Abstract
This survey addresses the critical challenge of deepfake detection amidst the rapid advancements in artificial intelligence. As AI-generated media, including video, audio and text, become more realistic, the risk of misuse to spread misinformation and commit identity fraud increases. Focused on face-centric deepfakes, this work traces the evolution from traditional single-modality methods to sophisticated multi-modal approaches that handle audio-visual and text-visual scenarios. We provide comprehensive taxonomies of detection techniques, discuss the evolution of generative methods from auto-encoders and GANs to diffusion models, and categorize these technologies by their unique attributes. To our knowledge, this is the first survey of its kind. We also explore the challenges of adapting detection methods to new generative models and enhancing the reliability and robustness of deepfake detectors, proposing directions for future research. This survey offers a detailed roadmap for researchers, supporting the development of technologies to counter the deceptive use of AI in media creation, particularly facial forgery. A curated list of all related papers can be found at https://github.com/qiqitao77/Awesome-Comprehensive-Deepfake-Detection.
Index Terms:
Generative Adversarial Network, Diffusion Model, Deepfake, Facial Forgery, Multi-modality
1 Introduction
Deepfake, encompassing video, audio, and text, has been transformed through advanced artificial intelligence technologies such as Variational Auto-Encoders (VAEs) [1], Generative Adversarial Networks (GANs) [2], and Diffusion Models (DMs) [3] to achieve unprecedented realism. Given that face-related deepfakes may have more severe societal and ethical ramifications than synthetic media of non-facial content, this work specifically focuses on facial deepfakes. It covers various facial manipulations such as face swapping, attribute editing, and face reenactment [4]. Those manipulations not only enhance personalized digital interactions on social platforms and mobile apps but also drive innovations in the entertainment industry, such as digitally resurrecting historical figures [^1] https://edition.cnn.com/2024/02/12/asia/suharto-deepfake-ai-scam-indonesia-election-hnk-intl/index.html or creating virtual actors  [^2] https://edition.cnn.com/2021/08/06/tech/tom-cruise-deepfake-tiktok-company/index.html.
Recent years have witnessed significant advances in deepfake generation, particularly with the emergence of diffusion models, which have revolutionized the creation of synthetic content. This evolution has greatly obscured the distinction between real and fake media, complicating the task for human observers and deepfake detectors. Furthermore, the scope of deepfake technology has expanded beyond single-modality (video or audio alone) to intricate multi-modal manipulations[5][6][7]. These manipulations blend fake audio, video, and text, introducing additional complexity to the detection process. Correspondingly, the deepfake detection mechanisms must evolve to address not only more sophisticated generation techniques but also the challenges posed by the integration of multiple media types.
Progress in deepfake detection over the past two years has addressed the challenges posed by these advanced generative methods and diverse modalities. Researchers have responded to the enhanced realism of synthetic content by innovating detection strategies at different stages of the detection pipeline, including the input level, feature level, and learning level. Some advances leverage new deep learning models, such as transformers [8] and vision-language models [9]. To address multi-modal deepfake detection, various fusion techniques, such as [10, 11, 12, 13], have been developed to utilize knowledge from different modalities effectively. These enhancements refine detection techniques to keep pace with the rapidly evolving quality and diversity of deepfakes, expanding efforts from traditional single-modal to comprehensive multi-modal detection strategies, and from passive to proactive solutions[14][15].
TABLE I: Comparison between this work and previous reviews. GAN: Generative Adversarial Network; AE: Auto-Encoder; DM: Diffusion Model.
Work Focused Tasks Generative Model Face-centric Multi-modal Released Date
Juefei-Xu et al.[16] Deepfake Generation and Detection GAN & VAE X Mar 2022
Pei et al.[4] Deepfake Generation and Detection GAN & VAE & DM X Mar 2024
Lin et al.[17] AIGC Detection LLM & DM X Feb 2024
Wang et al.[18] Deepfake Detection GAN X Nov 2023
Wang et al.[19] Reliability of Deepfake Detection GAN & VAE X Nov 2022
Ours Deepfake Detection GAN & VAE & DM Jun 2024
The rapid advancements and variations across different works in the field have motivated us to provide the most up-to-date survey, presented in this submission. As demonstrated in Table I, our survey is the first comprehensive review to specifically address facial forgery detection, ranging from single-modal to multi-modal methods. This unique focus distinguishes our survey from previous works in the field. For instance, while [17] explores detection across general fake samples, our survey specifically targets face-related generated media due to its significant potential for societal harm. In contrast to surveys such as [16][4], which allocate much of their discussion to deepfake generation, our work concentrates on detection techniques.
Unlike previous surveys that focused on either facial forgery [16, 4, 18, 19] or multi-modal deepfake [17, 20] separately, our work merges these two crucial areas, offering a holistic view of the current state and challenges of the field.
Our survey is exceptionally current, including the latest research up to June 2024, whereas most prior surveys only cover works released before 2023 [16, 18, 19, 20]. This ensures that we encompass the new emerging detection tasks besides conventional ones. Specifically, we include multi-modal deepfake which represents the trend in real-world scenarios, while existing works[16, 4, 18] merely discuss visually manipulated deepfakes. Besides, we also include the detection of deepfakes by novel generative techniques that are not considered in many previous reviews[16, 4, 18, 19], i.e., synthetic images generated by diffusion models. This survey incorporates the most advanced and state-of-the-art techniques in detection, providing readers with insights into cutting-edge methods and their effectiveness against evolving threats. As such, it serves as an invaluable resource for researchers and practitioners aiming to remain at the forefront of technological advancements in this rapidly evolving field.
More importantly, as illustrated in Figure 3, we offer a more detailed and systematic categorization of the literature than previously available, unlike previous studies  [16] and [4] that categorize deepfake detection methods into less than four types. We contend that our granular examination helps illustrate how different methods enhance detection capabilities in different aspects. By sorting the works according to the manipulation modalities they address, from single-modal to multi-modal detection, our survey clearly differentiates between the techniques used in various contexts.
Within the single-modal category, we distinguish between passive and proactive detection methods, categorizing them based on their objective—whether detecting deepfakes post-generation or preemptively preventing their creation. In passive detection, we recognize the generalization as a core issue, thus dividing the literature into basic and generalized detection. For each type, we further categorize works based on the type of knowledge utilized—artifacts or consistencies—and the stage of application, whether at the input level, model level, or learning level. We also delve into specialized methods designed for newer generative models like DMs, which have not been extensively covered in previous reviews. For proactive detection, we categorize the works based on the generation models they target, such as GANs or DMs.
As the first work comprehensively investigating the multi-modal facial forgery detection domain, predominantly focusing on audio-visual interactions, our survey explores various multi-modal knowledge utilization strategies, including independent learning, joint learning, matching-based approaches, etc.
In summary, our survey not only brings the academic community up to date with the latest developments but also introduces a more nuanced categorization and analysis of facial forgery detection techniques. This structured approach not only aids in understanding the complexities of the field but also inspires new research directions and methodological innovations, making it an essential resource for anyone involved in the study of this direction.
Scope of This Survey. This survey is structured as follows: Section 2 provides essential background information, defines key terminologies, discusses relevant datasets, and explains evaluation metrics for deepfake detection. Section 3 offers a detailed analysis of deepfake detection methods, categorizing them from single-modal to multi-modal and from passive to proactive approaches, as illustrated in Figure 3. Section 4 discusses current challenges in deepfake detection and suggests future research directions. The survey wraps up with Section 5, summarizing the main insights and contributions.
2 Background
2.1 Task Formulation
In this subsection, we formally define single-modal and multi-modal deepfake detection tasks discussed in our survey, demonstrating the types of input media and the forms of detection outputs under various scenarios.
Single-modal Detection. In single-modal cases, we denote a visual deepfake dataset as $\mathcal{D}=\{(\mathbf{X_{i}^{v}},Y_{i})\}_{i=1}^{N}$, where $\mathbf{X_{i}^{v}}\in\mathbb{R}^{T\times 3\times H\times W}$ representing a video with $T$ frames and $Y_{i}\in\{real,fake\}$ indicating the authenticity of $\mathbf{X_{i}^{v}}$. Given a detector $F_{\theta}:\mathbb{R}^{T\times 3\times H\times W}\mapsto\{real,fake\}$ parameterized with $\theta$, the aim is to optimize it in terms of the loss function, either simple cross-entropy loss for classification or advanced ones demonstrated in section 3.1.1.2.
$\theta=\operatorname*{arg\,min}\limits_{\theta^{\prime}}\mathbb{E}_{\mathcal{D}}L(F_{\theta^{\prime}}(\mathbf{X^{v}}),Y)$ (1)
Along with binary classification, spatial and temporal localization are proposed to facilitate detection. Spatial localization applies to image and video input, predicting a mask of the skeptically manipulated region, while temporal localization pertains exclusively to video inputs, requiring the output of a temporal mask indicating the segment(s) manipulated in the given video.
(a) Multi-modal audio-visual deepfake dataset FakeAVCeleb[5], figure modified from [5].
(b) Multi-modal visual-text deepfake dataset $DGM4$, figure from [6].
Figure 1: Multi-modal deepfake detection demonstration.
Multi-modal Audio-Visual Detection[10]. This task tackles the problem when the visual input is video and accompanied by audio stream input, either or both of which can be manipulated. In this case, we denote the paired audio-visual dataset in Figure 1(a) as $\mathcal{D}=\{(\mathbf{X_{i}^{v}},\mathbf{X_{i}^{a}},Y)\}_{i=1}^{N}$, where $\mathbf{X_{i}^{v}}$ is the video defined previously, $\mathbf{X_{i}^{a}}\in\mathbb{R}^{M\times c}$ represents the auditory waveform with $c$ channels and $M$ sample points, and $Y\in\{real,fake\}$ is the binary target label. Similar to the single-modal scenario, multi-modal audio-visual deepfake detection is to obtain an optimal detector $F_{\theta}:\mathbb{R}^{T\times 3\times H\times W}\times\mathbb{R}^{M\times c}\mapsto\{real,fake\}$ incorporating both visual and auditory inputs:
$\theta=\operatorname*{arg\,min}\limits_{\theta^{\prime}}\mathbb{E}_{\mathcal{D}}L(F_{\theta^{\prime}}(\mathbf{X^{v}},\mathbf{X^{a}}),Y)$ (2)
Optionally, a more fine-grained classification can also be performed, where $F_{\theta}$ further predicts $Y^{v}\in\{real,fake\}$ and $Y^{a}\in\{real,fake\}$ for each modality.
Multi-modal Text-Visual Detection and Grounding This task involves analyzing image-text pairs as illustrated in Figure 1(b), the dataset is denoted as $\mathcal{D}=\{(\mathbf{X_{i}^{v}},\mathbf{X_{i}^{t}},Y^{v}_{i},Y^{t}_{i},\mathbf{B^{v}_{i}},\mathbf{M^{t}_{i}},Y_{i})\}_{i=1}^{N}$, where $\mathbf{X^{v}_{i}}\in\mathbb{R}^{3\times H\times W}$ denotes the image input, $\mathbf{X^{t}_{i}}\in\mathbb{R}^{L\times d}$ represents text with $L$ tokens of $d$-dimension, $\mathbf{B^{v}_{i}}\in\mathbb{R}^{4}$ is the manipulated bounding box, $\mathbf{M^{t}_{i}}\in\mathbb{R}^{L\times 1}$ is manipulation mask of text modalities, $Y^{v}$, $Y^{t}$ and $Y$ indicate the authenticity of visual modality, text modality, and image-text pair respectively. Referring to Figure 1(b), $\mathbf{X^{v}}$ and $\mathbf{X^{t}}$ are the image and text in the middle, $\mathbf{B^{v}}$ and $\mathbf{M^{t}}$ correspond to the grounding annotation in the right part, $Y$, $Y^{v}$ and $Y^{t}$ are the binary authenticity labels in the left. In this task, the model needs to predict binary labels and manipulation masks, i.e., $\hat{Y}^{v},\hat{Y}^{t},\mathbf{\hat{B}^{v}},\mathbf{\hat{M}^{t}}=F_{\theta}(\mathbf{X^{v}},\mathbf{X^{t}})$. The detector is trained to minimize the designed loss given the dataset:
$\theta=\operatorname*{arg\,min}\limits_{\theta^{\prime}}\mathbb{E}_{\mathcal{D}}\\ L(F_{\theta^{\prime}}(\mathbf{X^{v}},\mathbf{X^{t}}),Y,Y^{v},Y^{t},\mathbf{B^{v}},\mathbf{M^{t}})$ (3)
2.2 Datasets for Facial Forgery Detection
In the rapidly evolving domain of deepfake detection, the formulation and assessment of detection algorithms heavily rely on various and exhaustive datasets. These datasets are broadly classified into single-modal and multi-modal categories based on the type of data they encompass.
Single-modal datasets are characterized by their focus on a singular form of evidence, providing specialized insights into detecting deepfakes through one sensory modality. Conversely, multi-modal datasets compile samples incorporating a variety of data types, e.g., visual, audio, and text. This integration mirrors the complex nature of deepfakes as they appear in real-world scenarios, offering a more holistic approach to detection by leveraging the richness of multiple data streams. We have summarized the database details related to facial forgery detection in Table 2.2.
TABLE II: Overview of Deepfake Datasets focusing on facial forgery. A: audio modality; V: visual modality; T: text; BC: binary classification; TG: temporal grounding; SG: spatial grounding; FS: face swap; FR: face reenactment; FE: face editing; TTS: text-to-speech; VC: voice conversion; WIG: whole image generation; TS: text swap.
Dataset* Year Type Tasks Manipulation Method #Subjects #Real #Fake
A V T
FaceForensics++ [21] 2019 Visual Deepfake by GAN/VAE BC - FS/FR - - $1,000$ $4,000$
DFD [22] 2019 BC - FS/FR - - $633$,06830683,0683 , 068
FaceShifter[23] 2020 BC - FS - - - $1,000$
DFDC [24] 2020 BC - FS - $960$ $23,654$ $104,500$
Celeb-DF[25] 2020 BC - FS - $59$ $905$,63956395,6395 , 639
DeeperForensics-1.0[26] 2020 BC - FS - $100$ $50,000$ $10,000$
WildDeepfake[27] 2020 BC - - - - $3,805$ $3,509$
KoDF[28] 2020 BC - FS - $403$ $62,166$ $175,776$
FFIW10K[29] 2021 BC/SG - FS - - $10,000$ $10,000$
ForgeryNet[30] 2021 BC/SG/TG - FS/FR - $5,400$ $99,630$ $121,617$
DF-Platter[31] 2023 BC - FS - $454$ $133,260$ $132,496$
DeepFakeFace[32] 2023 Visual Deepfake by DM BC - WIG - - $30,000$ $90,000$
DiffusionFace [33] 2024 BC - FS/WIG - - $30,000$ $600,000$
DiFF [34] 2024 BC - WIG/FS - $1,070$ $23,661$ $537,466$
FakeAVCeleb [5] 2021 Audio-Visual Deepfake BC SV2TTS FR - $500$ $500$ $19,500$
LAV-DF [35] 2022 BC/TG TTS FR - $153$ $36,431$ $99,873$
AV-Deepfake1M[7] 2023 BC/TG TTS FR - $2,068$ $286,721$ $860,039$
DGM4[6] 2023 Text-Visual Deepfake BC/SG - FS/FE TS - $77,426$ $152,574$
    Only publicly available datasets are listed.
2.2.1 Single-modal (visual) Deepfake Datasets
Initially, computer graphics techniques are favored to build up deepfake datasets. However, with the advancement of generative AI technologies, GANs, AEs, and DMs have increasingly become preferred methods for deepfake data generation [21, 25, 26, 24, 27, 28, 33, 32]. This shift towards generative AI techniques reflects the progress in the field, offering more sophisticated and convincing deepfakes by leveraging the power of AIGC methods to manipulate or generate content. We show some examples of visual-modal deepfake in Figure 2.
GANs/VAEs based Generated Datasets. The landscape of deepfake detection has been significantly enriched by datasets leveraging GANs and AEs. Pioneering datasets such as FaceForensics++ (FF++)[21], DeepFake Detection Challenge (DFDC) [24], Celeb-DF (and its successor Celeb-DF v2)[25], DeeperForensics-1.0[26], WildDeepfake[27], the Korean Deep-Fake Detection Dataset (KoDF)[28], FFIW10K[29], and ForgeryNet[30] stand at the forefront of this effort. These datasets utilize GANs or VAEs to create synthetic face images that spans various manipulations and scenarios.
FaceForensics++ (FF++)[21] sets a benchmark with its comprehensive collection of manipulations, including face swapping and reenactment through methods like DeepFakes, Face2Face, FaceSwap, and NeuralTextures. A subsequently proposed face swapping technique FaceShifter[23] is applied to the real videos in FF++[21], resulting in a new set of fake samples. DFD[22] enriches deepfake dataset by combining the manipulated and pristine videos recording paid and consenting actors. DFDC [24] challenges detection algorithms with a diverse set of forgery techniques, including eight different methods such as FSGAN [36] and StyleGAN[37], aimed at generating various facial forgeries.
Celeb-DF series [25] harnesses VAEs to craft deepfake images by replicating facial expressions across donor and target faces, pushing the envelope in visual manipulation realism. DeeperForensics-1.0[26] introduces the DeepFake Variational Auto-Encoder (DF-VAE), a learning-based face swapping framework to generate convincing forgeries. KoDF[28] diversifies the dataset pool by incorporating multiple synthesis models and focusing on Korean subjects, addressing the underrepresentation of Asian demographics in prior deepfake detection datasets.
FFIW10K[29] is a large-scale dataset specifically designed for multi-person face forgery detection, highlighting the complexity of detecting alterations among multiple genuine faces. DF-Platter[31] improves deepfake creation techniques using low-resolution videos and emphasizes a gender-balanced, ethnically diverse subject pool. WildDeepfake[27] collects samples from the internet, offering a real-world testbed for detection algorithms. To expand the detection challenge to include spatial and temporal localization tasks, ForgeryNet [30] introduces a comprehensive resource with more than $5,400$ subjects and $15$ manipulation techniques.
During the Covid-19 pandemic, Ko et al. [38] develop a new deepfake dataset featuring face masks by applying masks to images from well-known deepfake collections and noticed a significant drop in detection accuracy of current deepfake detection models when analyzing faces with masks. All these GANs/VAEs based deepfake datasets contribute uniquely to the realm of detecting forgery images, providing researchers with fundamental resources to address the challenges of deepfake detection.
DMs based Generated Datasets. Recent advancements in Diffusion Models have significantly enhanced the generation of visually compelling content. This has spurred the development of specialized datasets focusing particularly on facial forgery detection, such as DeepFakeFace [32], DiffusionFace [33], DiffusionDB-Face/JourneyDB-Face [39], and Diff[34]. It should be noted that most manipulations in the aforementioned DM-generated datasets are focused on the image level. Extending them to video-level deepfakes requires further exploration.
DeepFakeFace[32] employs three distinct DMs—Stable Diffusion v1.5, Stable Diffusion Inpainting, and InsightFace—each generating $30,000$ synthetic images to create a richly diverse dataset tailored for deepfake detection. DiffusionFace[33] stands as the inaugural dataset dedicated to diffusion-based facial forgery, covering a wide range of manipulations from unconditional generation to sophisticated face swapping techniques. It utilizes an impressive array of eleven diffusion models to achieve a new level of realism in forged facial imagery, showcasing the adaptability of diffusion technologies and establishing a benchmark for DMs-based deepfake detection. DiffusionDB-Face and JourneyDB-Face[39] introduce cutting-edge benchmarks for deepfake detection, emphasizing a variety of head poses, facial attributes, styles, and appearances.
The newly developed DiFF[34] is a substantial collection consisting of more than $500,000$ images synthesized using $13$ different generation methods. It utilizes $30,000$ carefully selected textual and visual prompts to ensure that the synthesized images maintain both high fidelity and semantic consistency.
While these datasets primarily support the detection of facial forgeries, the evolving landscape of deepfake content also necessitates broader detection capabilities as seen in datasets like RWDF-23 [40] and Zoom-DF [41]. The RWDF-23 dataset [40], launched in 2023, includes $2,000$ deepfake videos from four online media-sharing platforms, collected between January and May 2023. It encompasses videos in four languages, providing a diverse set of real-world deepfake content. Additionally, this dataset features viewer comments and interactions, and the authors provide a thorough analysis of deepfakes, covering aspects such as creators, manipulation strategies, intentions, and the context of real-world content production. Furthermore, the study in [41] introduces Zoom-DF tailored for remote meeting and video conferencing scenarios. Unlike traditional datasets that focus primarily on facial manipulation, Zoom-DF uniquely concentrates on altering participant movements, expanding the scope of deepfake detection challenges.
Refer to captionRefer to caption
Figure 2: Examples of single-modal deepfake datasets. Images in the top row are fake samples from FF++[21]. Images in the bottom row are generated by Stable Diffusion v1.5. Figure from [32].
2.2.2 Multi-modal Deepfake Datasets
Multi-modal deepfake datasets are becoming indispensable as they provide a more intricate testing environment that closely mirrors actual digital interactions, where deceptive content might include any combination of altered videos, synthesized speech, and manipulated text. The multi-modal deepfake datasets simulate the real-world scenarios of how manipulated content delivers misinformation more convincingly. The development of such datasets aims to refine detection algorithms by offering a broad spectrum of data types for analysis. As illustrated in Figure 1, either or both auditory and visual content are forged in audio-visual deepfake dataset, and text-visual deepfake dataset involves text-image pairs where each modality can be falsified.
Audio-Visual Deepfake Datasets Datasets like FakeAVCeleb [5] and Joint Audio-Video Deepfake [10] enhance the depth of deepfake detection research by including video deepfakes with corresponding synthesized, lip-synced audio tracks or by integrating audio alterations into existing visual datasets. These efforts, alongside others like DefakeAVMiT [42] and Trusted Media Challenge (TMC)[43], contribute substantially to building resources capable of addressing the nuanced challenges posed by audio-visual deepfakes. MMDFD[44] extends beyond audio-visual deepfake datasets by encompassing text modality represented as video subtitles, however, without releasing public access thus hindering its further comprehensive evaluation benchmarking.
Moreover, specialized datasets such as Localized Audio Visual DeepFake (LAV-DF)[35] and AV-Deepfake1M[7] focus on the precise detection of temporal forgeries and localization, targeting content manipulations that subtly change the sentiment. These datasets underscore the evolving sophistication in deepfake generation and the pressing need for equally sophisticated detection methodologies.
Text-Viusal Deepfake Datasets The DGM4 dataset[6] represents a significant advancement in deepfake detection, focusing specifically on facial forgery within multi-modal media contexts. Unlike other datasets that primarily address image-text mismatches, DGM4 challenges detectors to identify forgeries by analyzing both images and their associated text. This dataset does not only support binary classification but also involve the detailed task of marking manipulated sections within images and identifying forged text elements, enhancing its utility for complex multimedia forgery detection.
In general, these efforts[5, 43, 10, 35, 42, 7, 44, 6] encapsulate a significant step toward developing versatile resources for testing and improving detection methods across the complicate multi-modal deepfakes.
2.3 Evaluation
Evaluation Metrics In prior work involving only the visual modality, where deepfake detection is cast as a classification task, usually a binary classification task, the widely adopted evaluation metrics include $ACC$ (accuracy), $AUC$ (area under ROC curve), $F1-score$, $AP$(average precision), and $EER$ (equal error rate).
Under multi-modal scenarios, where the fine-grained detection for each modality is in a multi-label manner[12, 6], $mAP$ is also reported in addition to the aforementioned metrics.
In recent text-visual grounding works [6], $mIoU$, $IoU_{50}$, $IoU_{75}$ are used as evaluation metrics for visual grounding, where the bounding box of the manipulated region is predicted. Regarding text grounding, where token-level binary prediction is made to indicate the authenticity of each text token, the classification precision and recall are evaluated.
Evaluation Settings The prevalent evaluation for deepfake detection algorithms can be categorized into in-domain and cross-domain settings according to the distributional gap between the training and testing data.
In-domain evaluation refers to evaluating the detectors on testing data that follows the same distribution as training data. In prior works [21, 24, 26], practitioners conduct in-domain evaluation using non-overlapping testing and training data from the same dataset, or in a narrower sense, of the same manipulation type having identical forgery patterns. Under this setting, the evaluation results reflect the efficacy of detectors under a controlled conditions, providing insight into their performance in an idealized environment.
Cross-domain evaluation is the case that testing data and training data are of different distributions, requiring detectors to possess generalizability in real scenarios. The burgeoning generative techniques highlight the necessity of cross-domain evaluation to prevent detectors from overfitting on the specific manipulation types seen during training. There are three major settings to evaluate the cross-domain generalizability:
  • Cross-dataset: Training and testing sets are from different datasets, e.g., using FF++[21] for training and Celeb-DF[25] for testing.
  • Cross-manipulation: Training and testing occur on the same dataset but involve different manipulation types, e.g., training on FaceSwap manipulations from FF++ and testing on NeuralTextures. This setting requires detailed forgery type annotations besides real/fake labels.
  • Cross-postprocessing: Evaluates how detectors perform on deepfakes that have undergone various post-processing operations, e.g., JPEG compression or video compression rate change[26], simulating real-life scenarios where media might be compressed or otherwise altered before spreading. For example, detectors might be trained on high-quality, raw versions of FF++ and tested on a compressed, lower-quality version of the same dataset.
2.4 History and Challenges of Deepfake Datasets
The evolution of deepfake generation methods has closely paralleled advancements in generative AI technologies, transitioning from early techniques such as GANs and auto-encoders to the more recent and sophisticated Diffusion Models. This progression has necessitated a corresponding shift in deepfake detection strategies, moving from identifying simple visual artifacts to decoding subtler patterns that mirror genuine content, challenging both automated systems and human observers.
Moreover, the emergence of multi-modal deepfake datasets introduces new challenges, requiring detectors to understand and analyze the interactions across different data types, for example, video, audio and text. This integration reflects the real-world scenarios where deepfakes are used, adding layers of difficulty in detection tasks.
In the mean time, the manipulation techniques within deepfakes are also becoming more refined, targeting only specific faces of an image or select frames within a video sequence. This subtlety demands that deepfake detection systems not only identify the presence of fakes but also pinpoint the exact areas of manipulation with high precision.
These developments across various manipulation methods, modalities, and improved qualities highlight the ongoing need for adaptive detection methodologies in response to evolving deepfake datasets.
3 Method
In the following subsections, we review works predominantly published in the last three years, encompassing three major areas: single-modal detection methods, multi-modal detection methods, and the works related to trust-worthy deepfake detection. Given the extensive array of methods and the intricate technical details inherent in each category, it is impractical to cover all available techniques. Thus, in this survey, we will focus on highlighting key works within each category to delineate the evolutionary trajectory. Figure 3 illustrates our taxonomy of single- and multi-modal deepfake detection in detail.
{forest}
[…]
Figure 3: The taxonomy of deepfake detection methods.
Addition, tier=level5 [[11][152][155][156][157][35][13], tier=last] ] [Attention, tier=level5 [[13][158][159], tier=last] ] [MLP Mixer Layer, tier=level5 [[12], tier=last] ] [Multi-task Strategy, tier=level5 [[10][42][156][12][160],tier=last] ] [Regularization, tier=level5 [[158][159][42][157], tier=last] ] ] ] [Matching-based Learning, tier=level3 [[161][162], tier=last] ] [Others, tier=level3 [[163][164][165][166], tier=last] ] ] [Text-Visual, tier=level2 [[6][167][168][169], tier=last] ] ] ]
3.1 Single-modal Detection
In this section, we concentrate on works related to single-modal methods for facial forgery detection in images and videos. We categorize these methods into two major groups: passive detection methods and proactive detection methods.
3.1.1 Passive Detection Methods
Passive deepfake detection employs methods and techniques to identify manipulated videos or images based solely on the available content. This approach operates without active intervention or additional information from the creation process. It focuses exclusively on analyzing the existing content to extract artifacts and inconsistencies, without relying on external inputs or markers embedded during the content’s creation. In the following subsections, we review the works that target GAN/VAE generated deepfakes, covering basic and generalized detection methods. For both basic and generalized approaches, we further categorize previous works based on their distinct characteristics.
3.1.1.1   Basic Detection for GAN/VAE
Visual Artifacts-based Methods Early efforts in deepfake detection primarily focused on identifying visual artifacts. Afchar et al.[45] develop MesoNet, a streamlined model adept at analyzing mesoscopic features of images to detect forgeries. Rossler et al.[21] utilize the XceptionNet framework to discern subtle artifacts in manipulated images. In [46], Liu et al. develop Gram-Net to leverage global image characteristics to effectively differentiate between real and fake facial textures. Zhao et al.[47] introduce a multi-attentional framework that improves the detection of textural features indicative of deepfakes. Tan et al.[48] and their subsequent work [49] propose to focus on extract artifacts introduced by up-sampling operations, which are more subtle and often overlooked in traditional artifact analysis. It should be noted that recent methods, such as [47, 48, 49], have been experimentally validated for application in generalized deepfake detection.
Advancements continued with Corvi et al.[50] observing that GAN models produce noticeable artifacts during the process. Those artifacts are visible in the Fourier domain and exhibit unusual regular patterns in their auto-correlation. Chen et al. [51] propose using a single patch in conjunction with widely recognized Spatial Rich Model filters to extract artifacts for deepfake detection. Juet et al. [52] introduce the Global and Local Feature Fusion (GLFF) framework to strategically select informative patches for the detailed extraction of local artifacts. Zhang et al. [53] discover that real and synthetically generated images show significantly different latent Gaussian representations when subjected to an inverse diffusion process using a pre-trained diffusion model. By exploiting these distinctions, they can enhance and accentuate subtle artifacts within generated images, improving the effectiveness of their detection methods. Zhang et al. [54] introduce the Cross Appearance-Edge Learning detector, a tool specifically designed to recognize forgery by capturing both coarse and fine-grained appearance and edge patterns at multiple levels of granularity. This method effectively merges knowledge from both the appearance and edge domains to thoroughly expose a broad spectrum of forgery artifacts. Nguyen et al.[55] introduce a multi-task learning system with three key branches: classification, heatmap regression, and self-consistency regression. The heatmap regression and self-consistency regression branches are specially designed to pinpoint vulnerable areas within an image, areas that are particularly indicative of blending artifacts, thereby improving the framework’s ability to detect forgeries.
Consistencies-based Methods The field expanded into temporal artifact detection as researchers recognize the importance of dynamic inconsistencies in videos. Li et al.[56] utilize Long-term Recurrent Convolutional Networks to detect inconsistencies around blinking in videos, while Qi et al.[57] introduce DeepRhythm, which analyzes heartbeat rhythms to spot video forgeries. Haliassos et al.[58] design a spatio-temporal network to detect semantic inconsistencies in mouth movements within forgery videos. The study in [59] introduces a deepfake detection method called latent pattern detection, which identifies temporal inconsistencies across video frames. Haliassos et al.[170] develop a method that employs self-supervised learning to capture detailed video representations by using audio-visual correlations in genuine videos. The detector is then trained on both real and synthetic data to detect facial forgeries and to predict the learned representations from the initial stage for authentic videos. Pang et al.[60] propose MRE-Net consisted of two main components: Bipartite Group Sampling (BGS) and Multiple Rate Branch (MRB). BGS divides the video into bipartite groups to capture various face motion dynamics, while MRB captures short-term and long-term spatial-temporal inconsistencies, respectively. Wang et al.[61] capture spatio-temporal artifacts in one model based on disentangled learning strategy, dividing the convolutional weights of a 3D ResNet into spatial-based and temporal-based groups and updating the two groups alternatively during training. Xu et al.[62] introduce the Thumbnail Layout (TALL) strategy to enhance the ability to capture inconsistencies across frames in fake videos. Their following work TALL++[63] is built on top of TALL[62] by designing a graph reasoning block and introducing a consistency loss to better capture semantic-level inconsistency to improve generalizability. TI2Net [64] uses identity inconsistencies in video frames to authenticate clips. It encodes identity information from each frame into vectors and calculates the temporal identity embedding by analyzing the differences between these vectors. The temporal identity embedding effectively captures identity inconsistencies, which is crucial to determining the authenticity of the video input. In [65], Chen et al. focus on detecting compressed deepfake videos, which often present overlapping compression and tampering artifacts. In a recent work on video deepfake detection [171], Trevine et al. propose a two-stage training pipeline. The first stage uses self-supervised representation learning to enforce audio-visual correspondence with abundant real face videos, combining contrastive learning with a novel audio-visual complementary masking and feature fusion strategy. The second stage involves supervised downstream classification to train a classifier that takes advantage of the lack of cohesion in deepfake videos, thereby enhancing detection accuracy.
3.1.1.2   Generalized Detection for GAN/VAE
In the realm of deepfake detection, the capacity of generalized detection plays an important role in enhancing the adaptability of the detection methods against a wide range of manipulations. In this subsection, we delve into various categories of approaches such as input level, model level, and learning level methods. Each category provides their own perspectives and techniques to tackle the challenge of generalized deepfake detection.
Input level: Augmentation-based Methods Augmentation-based methods have emerged as a powerful strategy in deepfake detection by enriching training data with complex pseudo-fake examples. Wang et al. [66] develop an attention-based method focused on identifying key facial regions to lay the groundwork for targeted data enhancement. Specifically, the technique involves adaptively masking the facial areas most susceptible to manipulation, as identified through attention analysis. Building on the theme of using real image characteristics, Zhao et al. [67] introduce the Inconsistency Image Generator (I2G). This tool creates well-annotated training datasets by blending genuine images, thereby avoiding artifacts typically introduced by deep neural networks. Further exploring the diversity of forgery characteristics, Chen et al. [68] propose a method to improve the variety of forgeries the model encounters during training. By employing an adversarial training approach, this method not only increases the model’s exposure to diverse forgeries but also improves its ability to detect nuanced differences, making it more adept at identifying sophisticated fake content.
In further exploration, Shiohara et al. [69] utilize synthetic training data through ”self-blended images” (SBIs), which simulate common forgery indicators like blending seams and statistical discrepancies. Larue et al. [70] develop a data augmentation technique that subtly introduces soft discrepancies into synthetic face images. These slight adjustments help train the model, SeeABLE, to better recognize subtle inconsistencies. Li et al. [172] develop a method to enhance detection accuracy for low-quality face forgeries. They introduce a ”bleach generator” that synthesizes a corrective element that, when blended with the input forgery image, produces a ”bleached” image. This approach, bolstered by their image bleach generation technique and two specific loss constraints, effectively improves the detection results.
The prior methods create augmented forgery images through a pixel-level process. Recent works have proposed to conduct the augmentation in latent representation space [71] or feature space [72]. Yan et al. [71] introduce Latent Space Data Augmentation (LSDA), which innovatively manipulates the latent space to generate a broad spectrum of forgery variations. This approach minimizes the risk of overfitting to specific forgery methods and broadens the model’s ability to generalize across various types of manipulated content. In [72], FreqBlender advances pseudo-fake face generation by focusing on frequency knowledge. The approach is based on the hypothesis that faces consist of semantic, structural, and noise information, with structural information most related forgery traces. The authors employ a Frequency Parsing Network to adaptively segment critical frequency components, and create augmented fake samples by adaptively combining those components.
Input Level: Frequency-space based Methods Recently, researchers have expanded beyond RGB space analysis, developing sophisticated techniques that utilize frequency analysis to enhance deepfake detection. Qian et al.[73] highlight the role of frequency analysis in detecting forgery, introducing the Frequency in Face Forgery Network (F3-Net). This innovative two-stream collaborative learning framework leverages frequency-aware decomposed image components and local frequency statistics to identify forgery patterns effectively. Chen et al.[74] develop a face forgery detection method that examines connections between local image regions. They utilize the Multi-scale Patch Similarity Module (MPSM) to detect complex similarity patterns and the RGB-Frequency Attention Module (RFAM) to integrate multi-modal data. Li et al.[75] create an adaptive frequency feature generation module to dynamically identify forgery patterns in images. Their approach also introduces a single-center loss (SCL) technique to reduce intra-class variations among genuine faces and enhance class distinctions, improving feature discrimination in face forgery detection.
Building on these foundational works, subsequent research has introduced advanced approaches with additional complexities. Luo et al.[76] propose a method that leverages high-frequency knowledge through three modules: multi-scale high-frequency feature extraction, residual-guided spatial attention, and cross-modality attention, aimed at enhancing forgery detection by focusing on manipulation signs and harmonizing data aspects. Liu et al.[77] introduce Spatial-Phase Shallow Learning (SPSL), a method that utilizes up-sampling artifacts common in forged content. SPSL combines spatial analysis with phase spectrum insights, emphasizing local texture information to detect forgeries.
Further, Woo et al.[78] present the Attention-based Deepfake detection Distiller (ADD), which applies frequency domain learning and optimal transport theory in knowledge distillation. ADD uses frequency attention distillation to recover high-frequency components lost during compression and multi-view attention distillation to generate multiple attention vectors. Jeong et al.[79] create frequency-level perturbation maps to blur the distinctions between authentic and fabricated images. Their method, similar to adversarial learning, refines both the deepfake detector and the perturbation generator, enhancing the model’s ability to recognize forgeries from various GAN models, although adding complexity and computational demands. In [80], Wang et al. propose a Spatial-Frequency Dynamic Graph method using dynamic graph learning to analyze relation-aware features across spatial and frequency domains. Tan et al.[81] introduce FreqNet, a deepfake detection network that integrates frequency domain learning with a CNN classifier to significantly improve generalization across different scenarios. By applying convolutional layers to both phase and amplitude spectra, their approach detects a broad range of forgery indicators within the frequency domain. Doloriel et al.[82] find that frequency masking improves the generalization of deepfake detection across various generative models, achieving notable performance improvements.
Model Level: Transformer-Based Approaches Recently, transformers [8] have been employed in both static deepfake detection and temporal deepfake segmentation. Wang et al.[83] provide an overview (as of February 28, 2024) of fourteen works adopting Transformer architecture for deepfake detection.
The approach outlined in Saha et al. [84] utilizes a Transformer-based architecture for segmenting deepfake videos, specifically targeting frames that have been manipulated. This method integrates a Vision Transformer (ViT) to analyze spatial features using a Scaling and Shifting (SSF) strategy, alongside a Timeseries Transformer (TsT) that focuses on temporal aspects. The spatial data processed by the ViT is segmented into overlapping windows, which are then analyzed temporally by the TsT to detect deepfake manipulations effectively. In [85], Bai et al. introduce the Action-Units Relation Learning framework, which utilizes a newly designed Action Units Relation Transformer (ART) to enhance forgery detection by analyzing relationships between facial action units (AUs). ART incorporates both AU-specific and AU-agnostic branches to effectively uncover forgery clues. Furthermore, the framework includes the Tampered AU Prediction (TAP) technique, which manipulates AU-related regions to create challenging training samples, thereby refining the detection process.
Using advances in parameter-efficient tuning methods such as Adaptors[86, 88], Mixture of Experts[89, 88], and LoRA[89, 88], innovative strategies are being proposed to integrate these techniques into large pre-trained Vision-Transformers that serve as foundation models. These strategies elegantly address the data-demanding issue while adapting foundation ViTs to deepfake detection and circumvent distortion of originally learned rich knowledge during pre-training. The study by Nguyen et al. [90] investigates the effectiveness of self-supervised pre-trained vision transformers compared to supervised pre-trained transformers and traditional ConvNets. The findings suggest that transformers trained with self-supervised learning on broad datasets, not specifically tailored for deepfake detection, can generally outperform those trained through supervised methods. The DeepFake-Adapter [86] enhances a ViT by adding lightweight, dual-level adapter modules without altering the core architecture. In this method, Shao et al. propose to incorporate Globally-aware Bottleneck Adapters alongside the MLP layers to identify broad forgery cues and uses Locally-aware Spatial Adapters to interact with ViT features, enhancing the detection of localized anomalies in deepfake content. Similarly, Luo et al.[87] propose a novel Forgery-aware Adaptive Vision Transformer to boost generalization. They adapt ViT to deepfake detection by adding two specially designed components and keeping the vanilla ViT frozen to circumvent overfitting issue. The proposed framework is a dual-branch architecture, the CNN branch with Local-aware Forgery Injector aims to inject local information into ViT branch, and ViT-branch comprises vanilla ViT and Global-aware Forgery Adapter targets to learn generalizable features for deepfake detection. Liu et al.[89] propose the use of a mixture of low-rank experts to fully harness the potential of CLIP-ViT, thereby preserving existing knowledge while simultaneously expanding the model’s capacity for transferable deepfake detection. Kong et al. develop a novel method, MoE-FFD [88], which enhances face forgery detection by integrating LoRA and Adapter with the Vision Transformer. This approach effectively identifies forgery clues on both global and local scales. Additionally, their innovative Mixture-of-Experts modules boost the model’s capacity and adaptability, allowing for improved detection performance in a plug-and-play manner across different transformer architectures.
Model level: Vision-Language Models Researchers are harnessing the capabilities of pre-trained vision-language models (VLMs) [9], which are proving instrumental not only in universal deepfake detection, but also in enhancing reasoning processes.
Universal Detection: A primary focus has been developing methods that utilize the inherent capabilities of VLMs, such as CLIP [173], to evaluate their effectiveness across various scenarios. To enhance video content analysis, Han et al.[91] propose using the CLIP image encoder, paired with a newly designed side network-based decoder, to effectively extract both spatial and temporal cues from video clips. Unlike previous methodologies that directly utilized a frozen CLIP model, they update their CLIP-inspired model parameters by incorporating designed textual labels into the image augmentation process.
Reasoning: Recent advances have proposed using VLMs to incorporating reasoning to boost interpretability and context understanding in deepfake detection. For example, Tsigos et al. [92] propose a novel framework to evaluate how effectively explanation methods can interpret decisions made by deepfake detectors. This framework focuses on visual data and measures the accuracy of explanation methods by their ability to modify key image regions through adversarial attacks, assessing their impact on the detector’s verdict. The findings underscore the ability of this method to offer meaningful explanations based solely on visual input.
Building on the multimodal capabilities, Sun et al. develop Visual-Linguistic Face Forgery Detection (VLFFD)[93], which employs CLIP to improve face forgery detection. This approach enriches training data with detailed textual prompts and fine-grained images, employing a Coarse-and-Fine Co-training (C2F) framework that leverages both visual and linguistic knowledge. VLFFD improves forgery identification and provides in-depth interpretative insights into the locations and characteristics of the forgeries, markedly expressing its explanatory power. Zhang et al. [94] reformulate the deepfake problem as a Deepfake Detection Visual Question Answering (DD-VQA) task, which involves identifying image authenticity and providing explanations. They also create a new dataset and develop a Vision and Language Transformer-based framework that effectively detects deepfakes and generates high-quality explanations.
As Multimodal Large Language Models (MLLMs) have shown significant problem-solving capabilities in various vision fields, recent studies have begun exploring their application in deepfake detection. A recent study by [95] introduces the SHIELD benchmark to evaluate the effectiveness of MLLMs in face forgery detection across both visual and acoustic modalities. Tested under various conditions, including zero-shot and few-shot scenarios, MLLMs demonstrate superior capabilities in interpretability and multimodal reasoning in deepfake dtection. In parallel, Shan et al. [96] have validated the efficacy of MLLMs through extensive qualitative and quantitative experiments, highlighting the importance of effective prompt engineering to distinguish between real and AI-generated images effectively. In a recent work[97] Li et al. conduct extensive experiments to evaluate the performance of $13$ well-known language and multimodal models on the task of detecting fake images and explore how explicit reasoning impacts the accuracy of fake detection with these models.
Learning Level: Advanced Loss Functions While most deepfake detection methods traditionally rely on Softmax loss, Yin et al. [98] challenge the effectiveness of softmax loss for distinguishing subtle distinctions between real and fake images. They propose a method that utilizes both angular and Euclidean spaces to enhance class separability. By decoupling the magnitude and direction of embedding vectors, they introduce margin penalties in both domains to increase the separation between classes, thereby enhancing inter-class separability and intra-class compactness. Zhang et al. [99] further the development of discriminative features using a fractional center loss, designed to cluster real faces closer within the embedding space while distinctly separating them from fake ones. In their subsequent research, Zhang et al. [100] introduce a dual-level center loss approach that incorporates both patch-level and image-level center losses. This innovative method effectively reduces intra-class distances while expanding inter-class distances, significantly enhancing the discriminative capability of feature embeddings at both the micro (patch) and macro (image) levels.
Learning Level: Disentangled Representation Under the observation that representations learned by the detector contain both forgery and content information and thus remain recognizable to content clues like face identity and background, making the detector prone to overfit training data[101][102], disentanglement learning aims to decouple the learned features into task-relevant and task-irrelevant groups, while only the task-relevant group will be used for subsequent detection.
Liang et al.[101] initially explore a framework to disentangle the forgery-relevant and content-relevant features and remove content information by only employing the decoupled forgery feature for detection. The framework is devised with two encoders for content and forgery feature extraction, respectively. To ensure the effectiveness of disentanglement, a multi-task strategy and a contrastive constraint are proposed. The work of UCF[102] extends the framework by further decoupling the forgery-relevant features into manipulation-specific and manipulation-agnostic ones, targeting a more refined approach to feature disentanglement in deepfake analysis and achieving a better generalization ability across various forgery types. Lin et al.[103] also adopt this disentanglement framework with an additional focus on demographic-relevant features to improve the fairness of deepfake detection in addition to generalization ability.
Unlike the aforementioned works that decompose features from a high-level semantic perspective with the aid of multi-task strategy, Ba et al.[104] leverage a disentanglement module to extract multiple non-overlapping local features with the guidance of carefully derived Local Information Loss. These local features are subsequently fused and filtered out task-irrelevant information under the supervision of Global Information Loss.
Learning level: Reconstruction Researchers have innovated methods for identifying deepfakes by harnessing insights gained during the reconstruction of facial images. Khalid et al.[105] employ a One-class Variational Autoencoder (OC-VAE), trained solely on authentic images, to identify deepfakes by analyzing deviations in reconstruction scores. He et al.[106] utilize image re-synthesis techniques such as super-resolution and denoising to extract residual detection cues. Cao et al.[107] combine reconstruction with metric constraints and classification to distinguish real from forged images effectively. In MARLIN [108], Cai et al. use self-supervised learning to reconstruct accurate facial expressions from extensive online video data. These approaches leverage facial imagery reconstruction as a pivotal element for detection.
Learning level: Manipulation Localization The objective of manipulation localization in deepfake detection is to pinpoint altered pixels in fake samples. This approach not only reveals the specific changes made by attackers but also aids in discerning their motives, which is valuable for informing decisions made by authorities[109].
Li et al.[110] develop Face X-ray, a technique based on the assumption that face manipulations involve a blending process. Their method helps differentiate real from fake images by identifying the presence and location of blending boundaries, thereby assessing image authenticity. Dang et al.[111] emphasize the importance of detecting manipulated face images and precisely localizing altered areas. They employ an attention mechanism to refine feature map processing, which significantly enhances the ability to distinguish between genuine and counterfeit faces. Shuai et al.[112] introduce a model that effectively localizes forgery areas by utilizing RGB images and Spatial Rich Mode (SRM) noise residuals. Their method integrates multi-modal and multi-scale patch features to pinpoint potential forgery areas accurately. Dong et al.[116] argue that implicit identity leakage deteriorates detection generalization, and correspondingly propose an effective ID-unaware detection model, which is guided by authenticity and manipulation localization information. In [113], Zhang et al. create EditGuard, a system designed to detect manipulated regions within digital content. EditGuard embeds imperceptible watermarks that allow for the precise decoding of tampered areas. Recent works from Hu et al.[114] and Tantaru[115] have also contributed to this direction, with Tantaru developing a weakly-supervised method that enhances the localization of manipulated areas in images produced by deep manipulation techniques.
One significant challenge in manipulation localization is the high cost associated with annotating manipulated regions, which can be both time-consuming and expertise-intensive.
Learning level: Identity Discrepancy This approach shifts the focus from simply identifying if a face is real or fake to analyzing whether the facial identity extracted from the input maintains the subject’s biometric traits.
Cozzolino et al. [117] develop a system incorporating a facial feature extractor, a temporal network to identify biometric discrepancies, and a generative adversarial network that predicts subject-specific motions, enhancing the detection of inconsistencies in facial features. Nirkin et al. [118] analyze discrepancies between identity embeddings from the inner(target person) and outer (source person)facial regions. This comparison helps detect forgeries by highlighting inconsistencies in identity representations. Dong et al. [119] introduce the Identity Consistency Transformer, which uses identity vectors to discern differences between the inner and outer facial areas. Their model employs a consistency loss mechanism to align similar identities and differentiate dissimilar ones. Huang et al. [120] leverage implicit identity features of fake faces with two innovative loss functions: the Explicit Identity Contrast (EIC) loss, which attracts real samples to their true identities and repels fake ones, and the Implicit Identity Exploration (IIE) loss, a margin-based approach that ensures effective differentiation of genuine and manipulated faces by maintaining appropriate intra-class and inter-class distances.
This line of methods is believed to offer broad applicability across various types of manipulation, as it does not depend on the specific methods of manipulation.
Learning level: Knowledge Distillation Knowledge distillation techniques [174] have proven effective in enhancing the generalizability of deepfake detection, applicable both at the latent representation and/or prediction levels.
Woo’s Attention-based Deepfake Detection Distiller (ADD) [121] leverages high and low-quality forgery data to detect compressed deepfakes effectively. The innovative frequency attention distiller helps the student model recover lost high-frequency components, while the correlation distillation method improves pixel correlation preservation using the Sliced Wasserstein Distance metric [122].
Addressing scenarios without access to source domain data, Kim et al. introduce the Feature Representation Transfer Adaptation Learning (FReTAL) [123]. This method uses knowledge distillation to adapt detectors to new deepfake types efficiently, focusing on latent representations and prediction logits, assuming label availability for new forgeries. Similarly, the CoReD framework addresses sequential domain adaptation challenges by minimizing catastrophic forgetting within a teacher-student model setting [124]. This process ensures the knowledge preservation in both latent representation and prediction levels.
Furthermore, to improve deepfake detection across domains without needing label access for target samples, Lv et al.’s study introduces a bidirectional adaptation strategy[125]. Forward adaptation transfers detection capabilities to new forgeries via supervised training and adversarial feature adaptation, while backward adaptation uses self-distillation to further refine and maintain feature recognition and knowledge.
Learning level: Fine-grained Prediction Recent advancements highlight the importance of granularity and adaptability in facial forgery detection, refining the detection process to keep pace with the evolving complexity and variety of deepfake methods.
Guo et al.[126] approach the classification of forged images by representing the forgery attributes with multiple hierarchical labels. Their method classifies forgery from a broad category (fully-synthesized vs. partial-manipulated) down to specific forgery techniques. This hierarchical interpretation aids in better understanding the nuances of the forged image and enhances the separation of real and fake regions. By predicting the fine-grained and hierarchical path, their approach provides a more detailed analysis of the attributes that define forged images.
Guarnera et al. [127] take a similar hierarchical and multi-level approach to detect and recognize deepfakes. Their method begins by classifying images as real or AI-generated, then identifies the source of synthetic images—whether from GANs or DMs—and further pinpoints the specific architecture used.
Additionally, the controllable guide-space (GS) method introduced by Guo in [128] enhances feature discrimination across forgery domains. This method employs a well-designed guide-space that effectively segregates forgery domains and accentuates the distinction between real and forged domains in a finely controlled manner, improving the generalization of forgery features.
3.1.1.3   Detection for Diffusion Model
As diffusion models continue to evolve, their potential misuse for creating counterfeit content has raised significant concerns. The distinctness of their generation process necessitates specialized detectors to discern diffusion-generated images. Given that detection for diffusion models is in its nascent stages, most works are working on images from general domain, of which a subset contains facial forgery images. Nevertheless, these proposed frameworks can be applied to face-centric synthetic images without loss of generality.
The studies in [129] and [130] analyze outputs from diffusion models and concluded that, despite significant improvements in the quality of generated results, there are still identifiable clues that can distinguish these from real samples. Ricker et al. [129] extensively test whether current detection methods designed for GANs generated samples can identify images created by diffusion models. They find that while these methods initially struggled to tell DM-generated images apart from real ones, retraining them on DM images greatly improved their accuracy, even allowing them to detect images made by GANs. Corvi et al. [130] note that models like GLIDE and Stable Diffusion produce distinct spectral peaks, which can be effectively utilized to identify synthetic images. This spectral approach adds another dimension to the detection methodologies, harnessing unique signatures inherent in diffusion models.
Lorenz et al. [131] introduce the use of the lightweight multi Local Intrinsic Dimensionality (multiLID) technique, initially developed for detecting adversarial examples, to automatically detect synthetic images and identify the generator networks responsible for their creation. Following a different approach, Cheng et al. [34] present the Edge Graph Regularization (EGR) method, crafted to enhance the discriminative capabilities of detectors. This technique integrates edge graphs, derived via Sobel operators, as a regularization term into the empirical risk formulation. By doing so, it encourages the model to concurrently analyze the features of both the original images and the edge graphs, thereby refining its overall analytical precision.
Transitioning from intrinsic analysis to reconstruction accuracy, Wang et al. [132] introduce Diffusion Reconstruction Error (DIRE) that measures the discrepancy between an input image and its reconstructed version using a pre-trained diffusion model. Their method reveals that while images generated by diffusion models can be accurately reconstructed by the same type of models, real images do not replicate with the same fidelity. Expanding on the concept of reconstruction, Ma et al.[133] introduce the SeDID method, addressing the limitations of DIRE[132] by utilizing the intermediate steps in diffusion models. Unlike DIRE only focusing on the initial time-step information, SeDID captures critical information during the diffusion and reverse diffusion processes, potentially enhancing the accuracy of synthetic image detection. Building on this, Luo et al. [134] capitalize on conducting reconstructions within the latent space, enhancing efficiency in two key ways: it avoids the need for complete image inversion and reconstruction, and it reduces the risk of error accumulation associated with multiple inversion and reconstruction steps.
3.1.2 Proactive Detection Methods
Proactive detection methods for deepfakes utilize strategies that preemptively counteract deepfake manipulation before it occurs or facilitate the identification of manipulated content afterward. Unlike passive methods, which assess content post-creation, proactive approaches embed specific signals, such as adversarial perturbations, into the original content. These signals aid in verifying authenticity and integrity if the content is later altered. For effectiveness, these alterations visually resemble the original to evade detection while impairing the functionality of deepfake generation models. In the following subsections, we explore recent proactive detection efforts targeting deepfakes generated by both GANs and DMs.
3.1.2.1   Proactive Methods for GANs
Researchers have explored using adversarial attack techniques to inject perturbations into the original input, enhancing the resilience of images against unauthorized manipulations. Ruiz et al. [135] develop an approach using versatile adversarial attacks to prevent unauthorized facial modifications, improving the resilience of image translation systems through adversarial training for GANs. Similarly, Yeh et al. [136] use adversarial loss to increase the resistance of images to deepfake manipulations.
To further improve the resilience of perturbations, Huang et al. [137] introduce a poison perturbation generator that improves the effectiveness of perturbations using an alternating training strategy. Wang et al. [138] focus on improving the resilience of perturbations by targeting the Lab color space, particularly the decorrelated a and b channels. Expanding defensive strategies, Anejaet al. [139] develop a data-driven method that quickly produces tailored perturbations to redirect manipulation attempts toward a predefined outcome. Huang et al. [140] design the Cross-Model Universal Adversarial Watermark (CMUA-Watermark), broadening defense across various facial images and deepfake models. Asnani et al. [141] and [109] propose using real image templates to improve the detection accuracy of manipulation, helping differentiate between real and manipulated images.
Recent innovations have suggested conducting source tracing alongside injecting adversarial perturbations into the original input to enhance traceability against manipulations. Innovations by Wang et al. [142] and Wu et al. [143] underscore the importance of tracing original source images to boost detection system credibility. Yu et al. [14] enhance traceability by embedding traceable fingerprints directly into the generator’s parameters, linking outputs uniquely to their creators. Addressing gaps in proactive defenses, Tang et al. [144] introduce the Feature-Output ensemble UNiversal Disruptor (FOUND), which disrupts in the latent space and employs a gradient-ensemble algorithm to provide a universal attack pattern against multiple deepfake models, using minimal original-attribute images. In [145], Zhang et al. introduce Dual Defense, which embeds an invisible watermark in target faces to disrupt malicious face swapping while preserving traceability. Their approach utilizes a watermark embedding network that balances invisibility, adversariality, and traceability through perceptual adversarial encoding strategies. In a recent study, Wu et al. [146] demonstrate that traditional watermarking methods, when applied to AI-generated Deepfake content, can interfere with forgery detection. To address this issue, they propose ’AdvMark’, a plug-and-play solution that transforms robust watermarking systems into adversarial frameworks, retaining the ability to extract watermarks for provenance tracking.
3.1.2.2   Proactive Methods for DMs
Recent pioneering strategies in proactive deepfake detection for diffusion models, not specifically designed for facial forgery, include Cui et al.’s DiffusionShield [15], which enhances watermark uniformity, Fernandez et al.’s method [147] for embedding invisible watermarks in Latent Diffusion Models, and Wang et al.’s behavioral modification approach [148] to ensure traceability of protected content. These innovative methods establish foundational practices for applications in facial forgery detection.
Expanding the scope, Zhao et al. [149] examine the effectiveness of watermarking within diffusion models, with Cao et al. [150] extending proactive security measures to audio models. To enhance model transferability, Shim et al. [151] disrupt the latent encoding phase of deepfake generation using LEAT, which employs a Normalized Gradient Ensemble strategy to optimize disruptions across various models.
3.2 Multi-modal Deepfake Detection
The concurrence of fake content in multiple modalities leads to more convincing deepfakes compared to those visually manipulated ones. In such context, it becomes imperative to develop the capability to detect these intricate multi-modal deepfakes. In this subsection, we provide a comprehensive investigation of the more intricate and challenging multi-modal scenario in this subsection, where auditory or textual content can also be falsified together with the presence of visual manipulation.
3.2.1 Audio-Visual
Audio-visual deepfakes involve the manipulation of either or both the audio and visual streams in a video. The combination of manipulations on these two modalities potentially raises more serious ethical and societal concerns, e.g., manipulating politicians to make false claims convincingly by delivering a lip-synced video with fake audio.
The heterogeneous nature and the modality gap between audio and video make multi-modal audio-visual deepfake detection a more challenging task. Existing explorations on this task commonly design two-stream framework to incorporate multi-modal inputs. These efforts can be categorized into three main paradigms: independent learning, joint learning and matching-based learning.
(a) Independent learning for audio-visual deepfake detection.
(b) Joint learning for audio-visual deepfake detection.
(c) Matching-based learning for audio-visual deepfake detection.
Figure 4: Three main paradigms for multi-modal audio-visual deepfake detection.
Independent Learning The independent learning paradigm is a straightforward approach in audio-visual deepfake detection, where each modality is processed separately. This method then combines the results from each stream for a final decision [10][11][152][153].
As a baseline method presented in [10], two independent streams of CNN are proposed for audio and visual deepfake detection separately, and the final multi-modal detection result is obtained by combining the audio and visual predictions. In a preliminary attempt to evaluate the FakeAVCeleb[5] benchmark, Khalid et al.[11] ensemble predictions from independently learned uni-modal detectors. Furthermore, Hashmi et al.[152] devise an audio-visual network besides uni-modal networks and adopt ensemble learning to make the final decision based on predictions from audio, video, and audio-visual networks. Similarly, AVFakeNet[153] leverages Swin-Transformer [175] as backbones for both audio and visual branches for feature extraction and then integrates the result from both branches for the multi-modal prediction.
Despite its intuitive nature, the paradigm of independent learning might overlook the intrinsic relationships between audio and visual data, which can be critical in accurately detecting deepfakes.
Joint Learning This approach integrates audio and visual modalities to leverage multi-modal information for enhanced deepfake detection. The general network architecture of the joint learning paradigm encompasses two modal-specific branches taking as input corresponding modality content for feature extraction, i.e., video frames and audio waveform, and a fusion module to integrate features of different modalities for audio-visual detection. This paradigm mines the inter-modal relationship by feature fusion to enhance the multi-modal deepfake detection. Based on at which stage the modal-specific feature are fused together, joint learning-based detection algorithms can be further divided into intermediate and late fusion manners. Moreover, some existing late fusion-based audio-visual detectors leverage multi-task strategy or impose regularization to facilitate the learned multi-modal representations.
Intermediate Fusion: This method holds the assumption that simply combining the high-level abstract features extracted from both streams are insufficient to mine the inter-modal information. Zhou et al.[10] initially propose the joint audio-visual deepfake detection, handling cases in which either or both visual and auditory modalities have been manipulated. In their method, a sync-stream is designed to leverage the synchronization patterns in addition to separate branches for each modality, which fuses the intermediate features from distinct modalities by cross-attention mechanism. Yang et al.[42] employ a Transformer architecture, where audio and visual inputs go through a temporal-spatial encoder, and corresponding features of both modalities interact with each other via a decoder with bidirectional cross-attention module. Wang et al.[154] incorporate inspiration from how the biological system perceive the multi-modal environment, proposing visual and audio branches as different sensory systems, of which the obtained signals are then integrated by audio-visual attention mechanism acting like human multi-sensory neurons. Specifically, the intermediate audio and visual features attend to each other at all stages and are concatenated at the last few stages for further multi-modal detection.
Late Fusion: In this strategy, the audio and visual representations are only merged in the final stages of feature extraction. Techniques such as concatenation or addition are commonly used to merge modalities in late fusion [11][152][155][156][157][35][13]. Various advanced strategies, such as attention mechanism[158][159] and MLP mixer layer[12] are applied to enhance effective fusion.
While benchmarking the FakeAVCeleb[5] dataset, Khalid et al. [11] adapt multi-modal models from different tasks to audio-visual deepfake detection, in which the modal-specific features are integrated by concatenation. The audio-visual network presented in [152] uses a visual stream and an audio stream in parallel, taking as input the extracted video frames and mel-spectrogram of audio respectively, and the feature maps from two streams are then concatenated. DF-TransFusion[155] further explores this paradigm by using two branches to extract facial features by using self-attention for facial feature extraction and cross-attention for lip-audio synchronization, with a linear layer integrating the concatenated signals for final prediction.
Further, Liu et al. [157] propose a tri-branch framework to handle audio, video, and multi-scale frame inputs, where the extracted features are weighted based on inter-modal interactions and concatenated. Moreover, this method also designs Forgery Clues Magnification Transformer for feature extraction of each branch, where the tokens in the input sequence interact with each other to obtain the weighted sequence to magnify the intra-modal inconsistency, and Distribution Difference based Inconsistency Computing module to align information of audio and video modalities.
Dynamic weighted addition has proven effective in fusion strategies for multi-modal tasks. While addressing the temporal audio-visual deepfake localization task, Cai et al.[35] calculate the weighted sum of uni-modal boundary maps from corresponding branches as the fused boundary map, where the weights are dependent on uni-modal boundary maps themselves. Similarly, Wang et al.[13] leverage dynamic weighted averaging as late fusion approach in the proposed AVT2-DWF framework. Using a two-stream transformer architecture for uni-modal feature extraction, the fused information for multi-modal detection is obtained by weighted averaging the classification tokens in both uni-modal features. Notably, the weights AVT2-DWF are calculated by inter-modal attention mechanism.
More sophisticated than simple concatenation or addition, attention mechanisms and MLP mixer layers offer advanced alternatives for multi-modal feature fusion. Sree et al.[159] employ a multi-head self-attention module after concatenating features of different modalities. The attention mechanism further enhances the inter-modal interaction and benefits subsequent multi-modal detection. More than attention mechanism, Zou et al.[158] feed the concatenated audio-visual features into Transformer layers to fuse information from both modalities. Alternatively, Raza et al.[12] employ an MLP mixer layer that not only enhances intra-modal features but also facilitates the integration of inter-modal information by passing concatenated modal-specific features through an inter-modality MLP mixer layer.
Multi-task: While both intermediate and late fusion-based approaches focus on using fused information solely for multi-modal detection of audio-video pairs, there is a concern that they neglect the fine-grained differences between various types of audio-visual deepfakes. For example, scenarios where the video is real but the audio is fake, or vice versa, can confound feature extraction processes within each modality-specific branch[156]. In response, some works exploit a multi-task learning strategy to address the problem by simultaneously performing uni-modal and multi-modal detection[156, 10, 42, 12, 160].
The networks proposed in [10, 42] are tasked with three binary classification objectives: uni-modal audio and visual deepfake detection based on learned representations from corresponding branches and multi-modal detection based on the fused signals. Besides simultaneously introducing uni-modal detection tasks, Muppalla et al.[156] make the multi-modal detection more fine-grained, changing the original binary classification to a four-class one. Similarly, Raza et al.[12] reformulate multi-modal audio-visual deepfake detection as a multi-label classification task. They modify the classification head to incorporate fused signals and simultaneously perform classification tasks for both modalities. The multi-task strategy employed in these studies requires more fine-grained labels of each modality to provide supervision during training.
Unlike those incorporating fine-grained uni-modal detection tasks, Yu et al.[160] design an auxiliary task from a disparate perspective. Critiquing the reliance solely on audio-visual synchronization as insufficient for generalizable audio-visual deepfake detection, the paper emphasizes learning content-based cross-modal correlations. In this work, Yu et al. introduce a correlation distillation task, where well-trained audio and visual speech recognition models serve as teachers to obtain content-based cross-modal correlation, and the student detection network learns to model cross-modal correlation from teachers. This explicitly learned cross-modal correlation based on content helps to prevent overfitting.
Enhanced Representation: This line of work aims to learn more discriminative representations for effective detection during training. To this end, various strategies including regularization[42][157][158] and disentanglement[159] are employed.
Yang et al.[42] apply a contrastive loss and an additive angular margin loss to make the fused multi-modal features of manipulated and genuine samples more discriminative. Likewise, Zou et al.[158] also impose the contrastive constraint in cross-modal cases, thus the audio and visual features become more similar when they share common authenticity, i.e., both streams are real. Besides, Zou et al.[158] further introduce within-modal regularization to align uni-modal features with their respective targets, ensuring the uni-modal features are discriminative for corresponding uni-modal detection. The Distribution Difference based Inconsistency Computing module in [157] also leverages contrastive loss to impose regularization on the relationship between audio and video features based on Jensen-Shannon divergence.
In addition, Mis-AvoiDD[159] enhances representations for effective multi-modal deepfake detection by decoupling them into modal-invariant and modal-specific subspaces. Specifically, a modal-invariant encoder and two modal-specific encoders are devised to extract common and specific features from each modality. In modal-invariant subspace, the discrepancy between representations of different modalities is minimized to align modal-specific features. Furthermore, an orthogonal loss imposes constraints on the relationship between the modal-invariant and modal-specific features of the same modality as well as the relationship between the modal-specific features from different modalities. The orthogonal regularization ensures the distinctness between representation subspaces.
Matching-based Learning Distinct from fusion-based learning, matching-based learning leverages the natural synchronization found in genuine audio-visual content, which is typically absent in synthetic pairs[162, 161]. This approach focuses on matching audio and visual signals by evaluating the consistency between extracted features from each modality, rather than using traditional classification techniques. In this paradigm, networks employ modality-specific feature extractors and are trained on authentic content via contrastive learning, enhancing their generalization capabilities across various types of manipulation.
Cheng et al.[162] propose Voice-Face matching detection based on the observation verified in human perception and neurology field, which reveals the homogeneity between the face and voice of one person. The detection network involves only voice and face feature extractors, without fusion module and classification head. The predicted genuineness of the audio-visual pair is determined by the consistency of features from these two modalities. While training the feature extractors, contrastive InfoNCE[176] loss is applied to increase the similarity between matched voice and face feature pairs.
Tian et al.[161] adopt a matching-based Intra-Cross-Consistency method for multi-modal detection, where a matching score from the detection model is compared with a threshold to obtain detection results. Similar to [162], this approach also evaluates cross-consistency between audio and visual features to compute the matching score. The method addresses the trade-off observed in deepfake generation between temporal identity consistency and the synchronization of face motion with audio-visual elements by incorporating temporal identity intra-consistency. This method merely required self-supervised learning on real videos by a contrastive learning strategy, thus becoming generalizable.
Others There are several existing works on audio-visual deepfake detection leveraging disparate approaches and frameworks from the aforementioned paradigms.
Feng et al[165] handle the task from a novel perspective of anomaly detection, requiring training merely on real videos in a self-supervised manner. At the first stage of this method, an audio-visual network is trained to estimate the inter-modal synchronization score, which is used to extract heuristically designed features. Subsequently, an auto-regressive model is then proposed to learn the distribution of previously extracted feature sets. While performing detection, the samples with low likelihood estimated by the auto-regressive model are flagged as fake.
POI-Forensics[163] focuses on identity consistency instead of cross-modal inconsistency. Cozzolinno et al.[163] argue that any manipulation inevitably perturbs characteristics of the portrayed identity. Therefore, the authors utilize the indirect clue that the individual in the video is not consistent with the person who is claimed to be. While performing detection, this method requires a reference set of at least 10 videos of the person of interest, and the similarity index between the target video and reference videos would be used to determine prediction.
Haq et al.[164] introduce novel biological signal emotion psychology to improve deepfake detection. Based on the emotion extracted from each modality, the proposed neurosymbolic framework performs intra-modal emotion transition analysis and inter-modal reasoning. The former aims to explore the temporal emotion inconsistency, and the latter explores the cross-modal emotion consistency according to psychological theory. Notably, the proposed framework does not require training on deepfake datasets owing to the underlying psychology and emotion knowledge, leading to better generalizability.
Yu et al.[166] detect audio-visual deepfakes based on two augmented views of visual signals, proposing a two-stage framework PVASS-MDD to better capture the audio-visual inconsistency. The framework encompasses an auxiliary Predictive Visual-audio Alignment Self-supervision (PVASS) stage as well as the Multi-modal Deepfake Detection (MDD) stage. In the PVASS stage, the visual signals are augmented to two different views, each of which interacts with audio information via the Predictive VA Alignment module to obtain further corresponding VA representations, the self-supervised learning objective is to make the two VA representations similar. In the MDD stage, the two VA representations corresponding to two augmented visual views are concatenated to be fed into a classifier for multi-modal prediction. Besides cross-entropy loss, regularization terms including a contrastive loss and an auxiliary loss are additionally applied.
3.2.2 Text-Visual
Recent advances in text-visual deepfake detection have introduced sophisticated methods to address manipulations on text and visual modalities. Figure 5 illustrates the framework for this multi-modal detection and grounding task.
Figure 5: HAMMER[6] for text-visual deepfake detection and grounding. Figure from [6]
Shao et al. [167] introduce the DGM4 dataset, featuring samples manipulated through distinct image and text techniques. They propose the Hierarchical Multi-modal Manipulation Reasoning Transformer (HAMMER), which enhances cross-modal semantic alignment with a Manipulation-Aware Contrastive Loss, demonstrating its effectiveness on this comprehensive dataset.
Liu et al. [168] incorporate frequency domain analysis into RGB image evaluations using discrete wavelet transforms to capture forgery artifacts more effectively. Their approach integrates image and frequency data through a specialized encoder and a unified decoder, simplifying the optimization process while improving forgery detection.
Wang et al. [169] focus on modality-specific features for detecting and grounding multi-modal manipulations. To simultaneously preserve multi-modal alignment capability and exploit modality-specific features, the framework is devised with dual-branch cross-attention for bi-directional interaction between modalities and decoupled fine-grained classifiers, which perform modality-specific fine-grained manipulation classification task based on features of corresponding modality instead of leveraging an integrated token for both modalities in HAMMER[167]. Furthermore, an implicit manipulation query for each modality is proposed to facilitate reasoning for grounding using learnable queries.
3.3 Trustworthy Deepfake
Despite considerable advances in deepfake detection, from single-modal to multi-modal approaches, concerns persist regarding the robustness and trustworthiness of these detectors. For example, a recent study by Abdullah et al. [177] highlights significant vulnerabilities in current detection methods. The study indicates that attackers can easily tailor generative models to circumvent traditional defenses. Moreover, the misuse of widely deployed vision foundation models allows for the creation of adversarial deepfakes through subtle manipulations of image semantics, thereby undermining the effectiveness of established detection techniques. In the following subsections, we will explore recent developments in deepfake detection, focusing on adversarial attacks, backdoor attacks, and defensive strategies.
Adversarial Attack Adversarial attacks in deepfake detection aim to deceive manipulation detectors by subtly altering original images across various domains—pixel space, frequency domain, and latent space. Carlini et al. [178] highlight how adversarial strategies can significantly lower the accuracy of deepfake detectors. Supporting studies [179, 180, 181] confirm these vulnerabilities, emphasizing the need for robust detection mechanisms.
In the pixel domain, Neekhara et al. [182] explore how adversarial attacks target specific facial features, while Liao et al. [183] develop the Multi-Layers Key Semantic Region Selection method to target essential pixels. Zhu et al. [184] and Liu et al. [185] introduce less noticeable perturbations using frequency and spatial features to streamline the application across various images. In [186], Huang et al. introduce DeepNotch, a method to create realistic and undetectable DeepFakes by reducing artifact patterns without compromising image quality, using implicit spatial-domain notch filtering. For frequency-based attacks, Jia et al. [187] and Chen et al. [188] focus on disrupting frequency domain properties to fool detection systems, utilizing cross-task perturbed initialization to enhance attack efficacy. In [189], Wu et al. introduce TraceEvader, a novel evasion attack that manipulates high-frequency traces and introduces adversarial blur into low-frequency components. In latent space manipulations, Meng et al. [190], Shamshad et al. [191], and Li et al. [192] leverage the latent spaces of generative models to enhance the stealth and effectiveness of their attacks, merging semantic manipulations with adversarial perturbations to bypass forensic classifiers.
In a departure from traditional 2D adversarial attacks, Shahriyar’s studies [193] and the innovative ”AdvHeat” by Wang et al. [194] explore new dimensions such as video content dynamics and 3D adversarial face views.
Backdoor Attack Unlike adversarial attacks that manipulate input at runtime, backdoor attacks embed corrupted samples in training datasets. These samples appear legitimate, but contain Trojans that, when triggered, manipulate the model’s output to evade detection. Triggers can be placed in the pixel space [195, 196] or latent space [197], and can be selected from being manually specified [195] to stealthy [197, 196].
Sun et al. [195] propose to craft effective triggers by exploiting the vulnerabilities of deepfake detectors and select impactful samples with a Filtering-and-Updating Strategy. Han et al. [197] use an ”analysis-by-synthesis” approach to embed natural triggers in the latent space, involving Model Discrimination and Data Distribution to create independent poisoned samples.
Liang et al. [196] focus on embedding inconspicuous triggers, limiting perturbations to critical facial areas and employing a scalable trigger generator that creates transformation-sensitive patterns. This method enhances the stealth of triggers and addresses their sensitivity to image transformations, significantly improving the subtlety of backdoor attacks.
Discrepancy Minimization This approach focuses on reducing the detectable differences between genuine and manipulated images across various data representation domains, including pixel space, frequency space, and feature space.
In pixel space, Huang et al. [198] develop FakePolisher, which reduces artifacts by projecting images into a low-dimensional subspace. Their subsequent FakeRetouch [199] enhances image realism through strategic noise addition and image filtering. Similarly, Li et al. [200] and He et al. [201] improve facial data reconstruction by searching adversarial latent codes.
In the frequency domain, Dong et al. [202] balance spectral power distributions to make discrepancies less noticeable, while Lee et al. [203] introduce Spectrum Translation for refining image generation spectra.
For feature space discrepancies, Hou et al. [204] minimize differences by applying natural degradations in an adversarial manner. Ivanovska et al. [205] employ Denoising Diffusion Models to challenge detection systems, focusing on reconstructing images with specific denoising steps for robust testing. Additionally, Liu et al.’s TR-Net [206] efficiently removes intrinsic deepfake traces through coordinated generators and multiple discriminators, enhancing the method’s effectiveness across various detection systems.
Defense Strategies To enhance the resilience of deepfake detection systems against increasingly sophisticated attacks, researchers have developed robust defense strategies. Khan et al. introduce Adversarial Feature Similarity Learning (AFSL) [207], which improves detection robustness by optimizing feature similarity and implementing strategic regularizations. Additionally, Pinhasov et al. utilize Explainable Artificial Intelligence (XAI) to create interpretability maps, aiding in the robust classification of adversarial attacks [208]. The approach proposed in [209] improves the robustness of deepfake defenses through a destruction-aware constraint during training. [210] introduces FaceGuard that enhances defense against unseen adversarial attacks by generating various perturbations that evade detection, while concurrently learns to identify and remove these perturbations from images. Ensemble methods have also been prominent, with multiple researchers employing these strategies to capitalize on the diverse strengths of combined models, significantly enhancing detection accuracy and robustness against adversarial perturbations [211, 212, 213, 214]. These methods underscore a critical shift towards integrating advanced computational techniques to safeguard against the evolving landscape of digital forgery.
4 Challenges and Future Directions
As discussed in the last section, deepfake detection has achieved significant progress in recent years. However, various challenges persist due to the complexity of deepfake detection. In this section, we address these challenges and present potential avenues for future research.
Performance Generalization Across Scenarios Current deepfake detection methods show limited effectiveness in generalized deepfake detection, including across different datasets, manipulation methods, and post-processing techniques. The variations across these factors can result in significant statistical distribution differences between the training and testing sets. As pointed out in previous research on domain adaptation [215, 216] and domain generalization [217], ignoring these differences may lead to a dramatic performance drop when applying deepfake detectors in practical scenarios. Additionally, adversarial attacks pose another challenge on the robustness of models. Considering these challenges and the need for practical applicability, developing specific methods for generalized and robust deepfake detection becomes a promising and important direction.
Alignment Among Multi-modalities The transition from single-modal to multi-modal deepfake detection marks a critical shift in the field. Traditional detection techniques, which primarily focus on single modalities like video or audio, are increasingly inadequate as deepfake generation technologies evolve to incorporate multiple modalities, such as audio-visual or text-visual manipulations. This complexity requires advanced detection strategies that can handle the integration of various modalities. Recent research underscores the potential of exploiting correlations between different modalities to enhance detection accuracy. Future efforts should aim to further develop multi-modal detection frameworks, emphasizing improved synchronization and a deeper understanding of inter-modal dynamics.
Real-Time Analysis Deepfake detection, especially multi-modal detection, involves manipulation across different modalities and often occurs in the format of video/audio streams. In such cases, analysis and prediction can only be made after processing all streams from various modalities, which can be time-consuming. However, the potential risk introduced by the spread of deepfakes requires a real-time analysis and response [218]. Current methods focus more on prediction accuracy than efficiency, highlighting the need for further innovation and additional efforts to achieve real-time detection.
Expanding Deepfake Detection with LLMs A recent study [17] has explored using large language models (LLMs) for generation and detection. We believe that integrating LLMs into deepfake detection can be a promising direction, leveraging their powerful capabilities in detection and explainability. For example, LLMs can analyze subtleties in the text modality that may accompany deepfake videos, such as inconsistencies between spoken words and lip movements. Exploiting these discrepancies can enhance the analysis. Furthermore, LLMs can automatically generate explanations about detected deepfakes, detailing the reasons for the classification and providing evidence, which is crucial for trust and verification purposes.
5 Conclusion
This paper provides a comprehensive and up-to-date overview of deepfake detection, summarizing methods accepted by top-tier conferences and journals over the past three years. We first formally define the deepfake detection problem within our scope, followed by detailed datasets and evaluation metrics for deepfake detection, covering various modalities and generation models. Next, we systematically explore the mainstream deepfake detection methods and categorize them based on their characteristics. We believe that our careful categorization will help researchers quickly grasp the development trends in this field. Finally, we delve into the challenges and potential future directions in facial deepfake detection.