### v3

# Evolving from Single-modal to Multi-modal Facial Deepfake Detection: A Survey

Abstract
This survey addresses the critical challenge of facial deepfake detection amidst the rapid advancements in artificial intelligence. As AI-generated media, including video, audio, and text, become more realistic, the risk of misuse to spread misinformation and commit identity fraud increases. Focused on face-centric deepfakes, this work traces the evolution from traditional single-modality methods to sophisticated multi-modal approaches that handle audio-visual and text-visual scenarios. We provide comprehensive taxonomies of detection techniques, discuss the evolution of generative methods from auto-encoders and GANs to diffusion models, and categorize these technologies by their unique attributes. To our knowledge, this is the first survey of its kind. We also explore the challenges of adapting detection methods to new generative models and enhancing the reliability and robustness of deepfake detectors, proposing directions for future research. This survey offers a detailed roadmap for researchers, supporting the development of technologies to counter the deceptive use of AI in media creation, particularly facial forgery. A curated list of all related papers can be found at https://github.com/qiqitao77/Awesome-Comprehensive-Deepfake-Detection.
Index Terms:
Generative Adversarial Network, Diffusion Model, Deepfake, Facial Forgery, Multi-modality
1 Introduction
Deepfake, encompassing video, audio, and text, has been revolutionized by advanced artificial intelligence technologies such as Variational Auto-Encoders (VAEs) [1], Generative Adversarial Networks (GANs) [2], and Diffusion Models (DMs) [3], achieving unprecedented levels of realism. Given that face-related deepfakes may have more severe societal and ethical ramifications than synthetic media of non-facial content, this work specifically focuses on facial deepfakes. Facial deepfakes encompass various facial manipulations such as face swapping, attribute editing, and face reenactment [4]. While these manipulations can enhance personalized digital interactions on social platforms and drive innovations in the entertainment industry, such as digitally resurrecting historical figures [^1] https://edition.cnn.com/2024/02/12/asia/suharto-deepfake-ai-scam-indonesia-election-hnk-intl/index.html or creating virtual actors [^2] https://edition.cnn.com/2021/08/06/tech/tom-cruise-deepfake-tiktok-company/index.html, they also pose significant risks by potentially being used for malicious purposes [^3] https://edition.cnn.com/2024/02/04/asia/deepfake-cfo-scam-hong-kong-intl-hnk/index.html.
Recent years have witnessed significant advances in deepfake generation, including advanced generative methods and various modality manipulations, which is shown in Figure 1. The emergence of DMs, which have revolutionized the creation of synthetic content, has greatly obscured the distinction between real and fake media, complicating the task for human observers and deepfake detectors. Furthermore, the scope of deepfake technology has expanded beyond single-modality (video or audio alone) to intricate multi-modal manipulations[5][6][7]. The blending of fake audio, video, and text adds significant complexity to the detection process, increasing the challenges of identifying these manipulations effectively.
Figure 1: Advances in facial deepfake generation and corresponding detection in recent years.
To address these challenges, corresponding deepfake detection methods have been proposed. Combating with the emergent novel manipulations, researchers have developed various strategies and leveraged advanced larger deep learning models, such as transformers [8] and vision-language models [9], to enhance detection capabilities. To effectively utilize knowledge from different modalities, various fusion techniques [10, 11, 12, 13] have been developed, refining detection methods to keep pace with the rapidly evolving quality and diversity of deepfakes. Furthermore, detection efforts have expanded from passive to proactive solutions [14][15], highlighting the ongoing need to adapt and improve detection methods to counter increasingly sophisticated deepfake techniques.
TABLE I: Comparison between this work and previous reviews. GAN: Generative Adversarial Network; AE: Auto-Encoder; DM: Diffusion Model.
Work Focused Tasks Face-centric Multi-modal Generative Model Released Date
Juefei-Xu et al.[16] Deepfake Generation and Detection X GAN & VAE Mar 2022
Pei et al.[4] Deepfake Generation and Detection X GAN & VAE & DM Mar 2024
Lin et al.[17] AIGC Detection X LLM & DM Feb 2024
Wang et al.[18] Deepfake Detection X GAN Nov 2023
Wang et al.[19] Reliability of Deepfake Detection X GAN & VAE Nov 2022
Ours Facial Deepfake Detection GAN & VAE & DM Jul 2024
The rapid advancements and variations across different works in the field have motivated us to provide the most up-to-date survey presented in this submission. As demonstrated in Table I, our survey is the first comprehensive review to specifically address facial forgery detection, ranging from single-modal to multi-modal methods. This unique focus distinguishes our survey from previous works in the field. Specifically, this survey possesses following characteristics:
1) It particularly concentrates on facial deepfake detection. In contrast to earlier surveys such as [16][4], which allocate a large part of their discussion to deepfake generation, our work focuses on detection techniques due to the scarcity of comprehensive review for a myriad of existing detection methods. While [17] explores detection across general fake samples, our survey specifically targets face-related generated media due to its significant potential for societal harm.
2) It covers a wider range of facial deepfakes spanning diverse modalities. Specifically, we include multi-modal deepfake, which represents the trend in real-world scenarios, while existing works[16, 4, 18] about facial deepfake merely discuss visually manipulated deepfakes. This extended coverage allows us to provide a more comprehensive understanding of the advancements and issues in deepfake detection.
3) It is exceptionally current, including the latest research up to July 2024, whereas most prior surveys only cover works released before 2023 [16, 18, 19, 20]. This ensures we encompass the detection of fake content resulting from the novel generative techniques, which are not considered in many previous reviews[16, 4, 18, 19], such as those generated by DMs.
4) It offers a more detailed and systematic taxonomy of existing state-of-the-art facial deepfake detection techniques than previously available, as illustrated in Figure 3. Unlike studies such as [16] and [4], which categorize deepfake detection methods into fewer than four types, our granular examination illustrates how different methods enhance detection capabilities in various aspects.
In summary, our survey not only brings the community up-to-date with the latest developments, but also introduces a more nuanced categorization and analysis of facial forgery detection techniques. This structured approach not only aids in understanding the complexities of the field but also inspires new research directions and methodological innovations, making it an essential resource for anyone involved in the study of this direction.
Scope of This Survey. This survey is structured as follows: Section 2 provides essential background information, defines key terminologies, discusses relevant datasets, and explains evaluation metrics for deepfake detection. Section 3 offers a detailed analysis of deepfake detection methods, categorizing them from single-modal to multi-modal and from passive to proactive approaches, as illustrated in Figure 3. Section 4 discusses current challenges in deepfake detection and suggests future research directions. The survey wraps up with Section 5, summarizing the main insights and contributions.
2 Background
2.1 Task Formulation
In this subsection, we formally define single-modal and multi-modal deepfake detection tasks discussed in our survey, demonstrating the types of input media and the forms of detection outputs under various scenarios.
Single-modal Detection. In single-modal cases, we denote a visual deepfake dataset as $\mathcal{D}=\{(\mathbf{X_{i}^{v}},Y_{i})\}_{i=1}^{N}$, where $\mathbf{X_{i}^{v}}\in\mathbb{R}^{T\times 3\times H\times W}$ representing a video with $T$ frames and $Y_{i}\in\{real,fake\}$ indicating the authenticity of $\mathbf{X_{i}^{v}}$. Given a detector $F_{\theta}:\mathbb{R}^{T\times 3\times H\times W}\mapsto\{real,fake\}$ parameterized with $\theta$, the aim is to optimize it in terms of the loss function, either simple cross-entropy loss for classification or advanced ones demonstrated in section 3.1.1.2.
$\theta=\operatorname*{arg\,min}\limits_{\theta^{\prime}}\mathbb{E}_{\mathcal{D}}L(F_{\theta^{\prime}}(\mathbf{X^{v}}),Y)$ (1)
Along with binary classification, spatial and temporal localization are proposed to facilitate detection. Spatial localization applies to either image or video input, predicting a mask of the skeptically manipulated region, while temporal localization pertains exclusively to video inputs, requiring the output of a temporal mask indicating the segment(s) manipulated in the given video.
Multi-modal Audio-Visual Detection[10]. This task tackles the problem when the visual input is video and is accompanied by the audio stream input, either or both of which can be manipulated. In this case, we denote the paired audio-visual dataset in Figure 1(a) as $\mathcal{D}=\{(\mathbf{X_{i}^{v}},\mathbf{X_{i}^{a}},Y)\}_{i=1}^{N}$, where $\mathbf{X_{i}^{v}}$ is the video defined previously, $\mathbf{X_{i}^{a}}\in\mathbb{R}^{M\times c}$ represents the auditory waveform with $c$ channels and $M$ sample points, and $Y\in\{real,fake\}$ is the binary target label. Similar to the single-modal scenario, multi-modal audio-visual deepfake detection is to obtain an optimal detector $F_{\theta}:\mathbb{R}^{T\times 3\times H\times W}\times\mathbb{R}^{M\times c}\mapsto\{real,fake\}$ that incorporates both visual and auditory inputs:
$\theta=\operatorname*{arg\,min}\limits_{\theta^{\prime}}\mathbb{E}_{\mathcal{D}}L(F_{\theta^{\prime}}(\mathbf{X^{v}},\mathbf{X^{a}}),Y)$ (2)
Optionally, a more fine-grained classification can also be performed, where $F_{\theta}$ further predicts $Y^{v}\in\{real,fake\}$ and $Y^{a}\in\{real,fake\}$ for each modality.
Multi-modal Text-Visual Detection and Grounding This task involves analyzing image-text pairs as illustrated in Figure 1(b), the dataset is denoted as $\mathcal{D}=\{(\mathbf{X_{i}^{v}},\mathbf{X_{i}^{t}},Y^{v}_{i},Y^{t}_{i},\mathbf{B^{v}_{i}},\mathbf{M^{t}_{i}},Y_{i})\}_{i=1}^{N}$, where $\mathbf{X^{v}_{i}}\in\mathbb{R}^{3\times H\times W}$ denotes the image input, $\mathbf{X^{t}_{i}}\in\mathbb{R}^{L\times d}$ represents text with $L$ tokens of $d$-dimension, $\mathbf{B^{v}_{i}}\in\mathbb{R}^{4}$ is the manipulated bounding box, $\mathbf{M^{t}_{i}}\in\mathbb{R}^{L\times 1}$ is manipulation mask of text modalities, $Y^{v}$, $Y^{t}$ and $Y$ indicate the authenticity of visual modality, text modality, and image-text pair respectively. Referring to Figure 1(b), $\mathbf{X^{v}}$ and $\mathbf{X^{t}}$ are the image and text in the middle, $\mathbf{B^{v}}$ and $\mathbf{M^{t}}$ correspond to the grounding annotation in the right part, $Y$, $Y^{v}$ and $Y^{t}$ are the binary authenticity labels in the left. In this task, the model needs to predict binary labels and manipulation masks, i.e., $\hat{Y}^{v},\hat{Y}^{t},\mathbf{\hat{B}^{v}},\mathbf{\hat{M}^{t}}=F_{\theta}(\mathbf{X^{v}},\mathbf{X^{t}})$. The detector is trained to minimize the designed loss given the dataset:
$\theta=\operatorname*{arg\,min}\limits_{\theta^{\prime}}\mathbb{E}_{\mathcal{D}}\\ L(F_{\theta^{\prime}}(\mathbf{X^{v}},\mathbf{X^{t}}),Y,Y^{v},Y^{t},\mathbf{B^{v}},\mathbf{M^{t}})$ (3)
2.2 Datasets for Facial Forgery Detection
In the rapidly evolving domain of deepfake detection, the formulation and assessment of detection algorithms heavily rely on various and exhaustive datasets. These datasets are broadly classified into single-modal and multi-modal categories based on the type of data they encompass.
Single-modal datasets are characterized by their focus on a singular form of evidence, providing specialized insights into detecting deepfakes through one sensory modality. Conversely, multi-modal datasets compile samples incorporating a variety of data types, e.g., visual, audio, and text. This integration mirrors the complex nature of deepfakes as they appear in real-world scenarios, offering a more holistic approach to detection by leveraging the richness of multiple data streams. We have summarized the database details related to facial forgery detection in Table 2.2.
TABLE II: Overview of Deepfake Datasets focusing on facial forgery. A: audio modality; V: visual modality; T: text; BC: binary classification; TG: temporal grounding; SG: spatial grounding; FS: face swap; FR: face reenactment; FE: face editing; TTS: text-to-speech; VC: voice conversion; WIG: whole image generation; TS: text swap.
Dataset* Year Type Tasks Manipulation Method #Subjects #Real #Fake
A V T
FaceForensics++ [21] 2019 Visual Deepfake by GAN/VAE BC - FS/FR - - $1,000$ $4,000$
DFD [22] 2019 BC - FS/FR - - $633$,06830683,0683 , 068
FaceShifter[23] 2020 BC - FS - - - $1,000$
DFDC [24] 2020 BC - FS - $960$ $23,654$ $104,500$
Celeb-DF[25] 2020 BC - FS - $59$ $905$,63956395,6395 , 639
DeeperForensics-1.0[26] 2020 BC - FS - $100$ $50,000$ $10,000$
WildDeepfake[27] 2020 BC - - - - $3,805$ $3,509$
KoDF[28] 2020 BC - FS - $403$ $62,166$ $175,776$
FFIW10K[29] 2021 BC/SG - FS - - $10,000$ $10,000$
ForgeryNet[30] 2021 BC/SG/TG - FS/FR - $5,400$ $99,630$ $121,617$
DF-Platter[31] 2023 BC - FS - $454$ $133,260$ $132,496$
DeepFakeFace[32] 2023 Visual Deepfake by DM BC - WIG - - $30,000$ $90,000$
DiffusionFace [33] 2024 BC - FS/WIG - - $30,000$ $600,000$
DiFF [34] 2024 BC - WIG/FS - $1,070$ $23,661$ $537,466$
FakeAVCeleb [5] 2021 Audio-Visual Deepfake BC SV2TTS FR - $500$ $500$ $19,500$
LAV-DF [35] 2022 BC/TG TTS FR - $153$ $36,431$ $99,873$
AV-Deepfake1M[7] 2023 BC/TG TTS FR - $2,068$ $286,721$ $860,039$
PolyGlotFake[36] 2024 BC TTS/VC FR - - $766$ $14,472$
DGM4[6] 2023 Text-Visual Deepfake BC/SG - FS/FE TS - $77,426$ $152,574$
    Only publicly available datasets are listed.
2.2.1 Single-modal (visual) Deepfake Datasets
Initially, computer graphics techniques are favored to build up deepfake datasets. However, with the advancement of generative AI technologies, GAN, VAE, and DM have become increasingly preferred methods for deepfake data generation [21, 25, 26, 24, 27, 28, 33, 32]. This shift towards generative AI techniques reflects the progress in the field, offering more sophisticated and convincing deepfakes by leveraging the power of AIGC methods to manipulate or generate content.
GANs/VAEs based Generated Datasets. The landscape of deepfake detection has been significantly enriched by datasets leveraging GANs and VAEs. Pioneering datasets such as FaceForensics++ (FF++)[21], DeepFake Detection Challenge (DFDC) [24], Celeb-DF (and its successor Celeb-DF v2)[25], DeeperForensics-1.0[26], WildDeepfake[27], the Korean Deep-Fake Detection Dataset (KoDF)[28], FFIW10K[29], and ForgeryNet[30] stand at the forefront of this effort. These datasets utilize GANs or VAEs to create synthetic face images that span various manipulations.
FaceForensics++ (FF++)[21] sets a benchmark with its comprehensive collection of manipulations, including face swapping and reenactment through methods like DeepFakes, Face2Face, FaceSwap, and NeuralTextures. A subsequently proposed face swapping technique FaceShifter[23] is applied to the real videos in FF++[21], resulting in a new set of fake samples. DFD[22] enriches the deepfake dataset by combining manipulated and pristine videos that record paid and consenting actors. DFDC [24] challenges detection algorithms with a diverse set of forgery techniques, including eight different methods such as FSGAN [37] and StyleGAN[38], aimed at generating various facial forgeries.
Celeb-DF series [25] harnesses VAEs to craft deepfake images by replicating facial expressions across donor and target faces, pushing the envelope in visual manipulation realism. DeeperForensics-1.0[26] introduces the DeepFake Variational Auto-Encoder (DF-VAE), a learning-based face swapping framework to generate convincing forgeries. KoDF[28] diversifies the dataset pool by incorporating multiple synthesis models and focusing on Korean subjects, addressing the underrepresentation of Asian demographics in previous deepfake detection datasets.
FFIW10K[29] is a large-scale dataset specifically designed for multi-person face forgery detection, highlighting the complexity of detecting alterations among multiple genuine faces. DF-Platter[31] improves deepfake creation techniques using low-resolution videos and emphasizes a gender-balanced, ethnically diverse subject pool. WildDeepfake[27] collects samples from the internet, offering a real-world testbed for detection algorithms. To expand the detection challenge to include spatial and temporal localization tasks, ForgeryNet [30] introduces more than $5,400$ subjects and $15$ manipulation techniques. During the Covid-19 pandemic, Ko et al. [39] develop a new deepfake dataset featuring face masks by applying masks to images from well-known deepfake collections and noticed a significant drop in detection accuracy of current deepfake detection models when analyzing faces with masks.
DMs based Generated Datasets. Recent advancements in Diffusion Models have significantly enhanced the generation of visually compelling content. This has spurred the development of specialized datasets focusing particularly on facial forgery detection, such as DeepFakeFace [32], DiffusionFace [33], DiffusionDB-Face/JourneyDB-Face [40], and Diff[34]. It should be noted that most manipulations in the aforementioned DM-generated datasets are focused on the image level. Extending them to video-level deepfakes requires further exploration.
DeepFakeFace[32] employs $3$ distinct DMs—Stable Diffusion v1.5, Stable Diffusion Inpainting, and InsightFace—each generating $30,000$ synthetic images to create a richly diverse dataset tailored for deepfake detection. DiffusionFace[33] is the inaugural dataset dedicated to diffusion-based facial forgery, covering a wide range of manipulations from unconditional generation to sophisticated face swapping techniques. It utilizes $11$ diffusion models to achieve a new level of realism in forged facial imagery, showcasing the adaptability of diffusion technologies and establishing a benchmark for DMs-based deepfake detection. DiffusionDB-Face and JourneyDB-Face[40] introduce cutting-edge benchmarks for deepfake detection, emphasizing a variety of head poses, facial attributes, styles, and appearances. The newly developed DiFF[34] is a substantial collection consisting of more than $500,000$ images synthesized using $13$ different generation methods. It utilizes $30,000$ carefully selected textual and visual prompts to ensure that the synthesized images maintain both high fidelity and semantic consistency.
Recently, Yan et al. create a large-scale and diverse dataset called DF40 [41], featuring $40$ distinct deepfake techniques, including GANs-based and DMs-based methods.. Using DF40, they conduct comprehensive evaluations with 4 standard protocols and $7$ representative detection methods, resulting in more than $2,000$ evaluations. This extensive analysis yields several insightful findings that contribute significantly to the field. For example, CLIP-large emerged as the most powerful baseline, significantly outperforming Xception and demonstrating the ability to generalize to non-face AIGCs even when trained solely on face data.
Although these datasets primarily support the detection of facial forgeries, the evolving landscape of deepfake content also requires broader detection capabilities, as seen in data sets such as RWDF-23 [42], Zoom-DF [43], and AI-Face dataset [44]. The RWDF-23 dataset [42], launched in 2023, includes $2,000$ deepfake videos from four online media-sharing platforms, collected between January and May 2023. It encompasses videos in four languages, providing a diverse set of real-world deepfake content. Additionally, this dataset features viewer comments and interactions, and the authors provide a thorough analysis of deepfakes, covering aspects such as creators, manipulation strategies, intentions, and the context of real-world content production. Furthermore, the study in [43] introduces Zoom-DF designed for remote meetings and video conferencing scenarios. Zoom-DF uniquely concentrates on altering participant movements, expanding the scope of deepfake detection challenges. To assess the fairness of facial deepfake detectors, Li et al. introduce the AI-Face dataset [44]. This groundbreaking dataset is the first of its kind to reach a million-scale and includes demographically annotated AI-generated face images.
(a) Multi-modal audio-visual deepfake dataset FakeAVCeleb[5], figure modified from [5].
(b) Multi-modal visual-text deepfake dataset $DGM4$, figure from [6].
Figure 2: Multi-modal deepfake detection.
2.2.2 Multi-modal Deepfake Datasets
Multi-modal deepfake datasets are becoming indispensable as they provide a more intricate testing environment that closely mirrors actual digital interactions, where deceptive content might include any combination of altered videos, synthesized speech, and manipulated text. The multi-modal deepfake datasets simulate real-world scenarios of how manipulated content delivers misinformation more convincingly. The development of such datasets aims to refine detection algorithms by offering a broad spectrum of data types for analysis. As illustrated in Figure 2, either or both auditory and visual content is forged in audio-visual deepfake datasets, and text-visual deepfake datasets involve text-image pairs where either modality can be falsified.
Audio-Visual Deepfake Datasets Datasets such as FakeAVCeleb [5] and Joint Audio-Video Deepfake [10] improve the depth of deepfake detection research by including video deepfakes with corresponding synthesized, lip-synced audio tracks or by integrating audio alterations into existing visual datasets. These efforts, along with others such as DefakeAVMiT [45] and Trusted Media Challenge (TMC)[46], contribute substantially to building resources capable of addressing the nuanced challenges posed by audio-visual deepfakes. MMDFD[47] extends beyond audio-visual deepfake datasets by encompassing the text modality represented as video subtitles, however, without releasing public access thus hindering its further comprehensive evaluation benchmarking. PolyGlotFake[36] enriches the language diversity of the audio-visual deepfakes, characterized as the first multi-lingual deepfake dataset covering $7$ languages in acoustic content.
Moreover, specialized datasets such as Localized Audio Visual DeepFake (LAV-DF)[35] and AV-Deepfake1M[7] focus on the precise detection of temporal forgeries and localization, targeting content manipulations that subtly change the sentiment. These datasets underscore the evolving sophistication in deepfake generation and the pressing need for equally sophisticated detection methodologies.
Text-Viusal Deepfake Datasets The DGM4 dataset[6] represents a significant advancement in deepfake detection, focusing specifically on facial forgery within multi-modal media contexts. Unlike other datasets that primarily address image-text mismatches, DGM4 challenges detectors to identify forgeries by analyzing both images and their associated text. This dataset does not only support binary classification but also involve the detailed task of marking manipulated sections within images and identifying forged text elements, enhancing its utility for complex multimedia forgery detection.
In general, these aforementioned efforts, such as [5, 46, 10, 35, 45, 7, 47, 6], encapsulate a significant step toward developing versatile resources for testing and improving detection methods across the complicate multi-modal deepfakes.
2.3 Evaluation
Evaluation Metrics In prior work involving only the visual modality, where deepfake detection is usually cast as a binary classification task, the widely adopted evaluation metrics include $ACC$ (accuracy), $AUC$ (area under ROC curve), $F1-score$, $AP$(average precision), and $EER$ (equal error rate). Under multi-modal scenarios, where the fine-grained detection for each modality is in a multi-label manner[12, 6], $mAP$ (mean Average Precision) is also reported in addition to the aforementioned metrics.
In recent text-visual grounding works [6], $mIoU$ (mean Intersection over Union), $IoU_{50}$, $IoU_{75}$ are used as evaluation metrics for visual grounding, where the bounding box of the manipulated region is predicted. Regarding text grounding, where a token-level binary prediction is made to indicate the authenticity of each text token, the classification precision and recall are evaluated.
Evaluation Settings The prevalent evaluation for deepfake detection algorithms can be categorized into in-domain and cross-domain settings according to the distributional gap between the training and testing data.
In-domain evaluation refers to evaluating the detectors on the testing data following the same distribution as the training data. In previous work such as [21, 24, 26], practitioners perform an in-domain evaluation using non-overlapping testing and training data from the same dataset, or in a narrower sense, of the same manipulation type having identical forgery patterns. Under this setting, the evaluation results reflect the efficacy of detectors under a controlled conditions, providing insight into their performance in an idealized environment.
Cross-domain evaluation refers to the scenario where the test data and training data come from different distributions, requiring detectors to have generalizability in real-world situations. The evolving generative techniques highlight the necessity of cross-domain evaluation to prevent detectors from overfitting on the specific manipulation types seen during training. There are three major settings to evaluate the cross-domain generalizability:
  • Cross-dataset: Training and testing sets are from different datasets, e.g., using FF++[21] for training and Celeb-DF[25] for testing.
  • Cross-manipulation: Training and testing occur on the same dataset but involve different manipulation types, e.g., training on FaceSwap manipulations from FF++ and testing on NeuralTextures also from FF++. This setting requires detailed forgery type annotations besides real/fake labels.
  • Cross-postprocessing: Evaluates how detectors perform on deepfakes that have undergone various post-processing operations, e.g., JPEG compression or video compression rate change[26], simulating real-life scenarios where media might be compressed or otherwise altered before spreading. For example, detectors might be trained on high-quality, raw versions of FF++ and tested on a compressed, lower-quality version of the same dataset.
2.4 History and Challenges of Deepfake Datasets
The evolution of deepfake generation methods has closely paralleled advancements in generative AI technologies, transitioning from early techniques such as GAN/VAE[21, 24, 22, 25] to the more recent and sophisticated DM[32, 33, 34]. This progression has necessitated a corresponding shift in deepfake detection strategies, moving from identifying simple visual artifacts to decoding subtler patterns that mirror genuine content, challenging both automated deepfake detectors and human observers.
Moreover, the emergence of multi-modal deepfake datasets[5, 35, 7, 36, 6] introduces new challenges, requiring detectors to understand interactions across different data types, for example, video, audio, and text. This integration reflects the real-world scenarios where deepfakes are used, adding layers of difficulty in detection tasks.
In the mean time, the manipulation techniques within deepfakes are also becoming more refined, targeting only specific facial regions of an image[29, 30, 6] or selecting frames within a video sequence[35, 7]. This subtlety demands that detection systems not only identify the presence of fakes, but also pinpoint the exact areas of manipulation.
In summary, these developments across various manipulation methods , modalities, and improved qualities highlight the ongoing need for adaptive detection methodologies in response to evolving deepfake datasets.
3 Method
In the following subsections, we review the works published predominantly in the last three years, encompassing the following major areas: single-modal detection methods, multi-modal detection methods. Given the extensive array of methods and the intricate technical details inherent in each category, it is impractical to cover all available techniques. Thus, in this survey, we will focus on highlighting key works within each category to delineate the evolutionary trajectory. Figure 3 illustrates our taxonomy of single- and multi-modal deepfake detection in detail.
{forest}
[…]
Figure 3: The taxonomy of facial deepfake detection methods.
Addition, tier=level5 [[11][169][173][174][175][35][13], tier=last] ] [Attention, tier=level5 [[13][176][177], tier=last] ] [MLP Mixer Layer, tier=level5 [[12], tier=last] ] [Multi-task Strategy, tier=level5 [[10][45][174][12][178],tier=last] ] [Regularization, tier=level5 [[176][177][45][175], tier=last] ] ] ] [Matching-based Learning, tier=level3 [[179][180][181], tier=last] ] [Others, tier=level3 [[182][183][184][185][186], tier=last] ] ] [Text-Visual, tier=level2 [[6][187][188][189], tier=last] ] ] ]
3.1 Single-modal Detection
In this section, we concentrate on works related to single-modal methods for facial forgery detection in images and videos. We categorize these methods into two major groups: passive detection methods and proactive detection methods.
3.1.1 Passive Detection Methods
Passive deepfake detection employs methods and techniques to identify manipulated videos or images based solely on the available content. This approach operates without active intervention or additional information from the creation process. It focuses exclusively on analyzing the existing content to extract artifacts and inconsistencies, without relying on external inputs or markers embedded during the content’s creation. In the following subsections, we first review the works that target GAN/VAE generated deepfakes, covering basic and generalized detection methods. For both basic and generalized approaches, we further categorize previous works based on their distinct characteristics. Then, we review the works about detecting DMs-generated deepfakes and sequential deepfake detection.
3.1.1.1   Basic Detection for GAN/VAE
In deepfake detection, artifacts usually refer to unintended irregularities that appear when creating or manipulating digital media using AI technologies. These imperfections can include mismatches in lighting and color [48], unusual facial textures [49], frequency domains [53], inconsistent edge blending [58], among others, that are not typically present in genuine media. In addition, inconsistencies can manifest as temporal anomalies in video sequences, where facial movements [62], identities [71], or noise patterns [72, 73], among others, do not maintain continuity between frames. Detecting these artifacts and inconsistencies is crucial to identifying deepfakes, as they serve as important indicators that the input media has been altered or synthetically generated.
Visual Artifacts-based Methods Early efforts in deepfake detection primarily focused on identifying visual artifacts. Afchar et al.[48] develop MesoNet, a streamlined model adept at analyzing mesoscopic features of images to detect forgeries. Rossler et al.[21] utilize the XceptionNet framework to discern subtle artifacts in manipulated images. In [49], Liu et al. develop Gram-Net to leverage global image characteristics to effectively differentiate between real and fake facial textures. Zhao et al.[50] introduce a multi-attentional framework that improves the detection of textural features indicative of deepfakes. Tan et al.[51] and their subsequent work [52] propose to focus on extract artifacts introduced by up-sampling operations, which are more subtle and often overlooked in traditional artifact analysis. It should be noted that recent methods, such as [50, 51, 52], have been experimentally validated for application in generalized deepfake detection.
Advancements continued with Corvi et al.[53] observing that GAN models produce noticeable artifacts during the process. Those artifacts are visible in the Fourier domain and exhibit unusual regular patterns in their auto-correlation. Chen et al. [54] propose using a single patch in conjunction with widely recognized Spatial Rich Model filters to extract artifacts for deepfake detection. Juet et al. [55] introduce the Global and Local Feature Fusion (GLFF) framework to strategically select informative patches for the detailed extraction of local artifacts. Zhang et al. [56] discover that real and synthetically generated images show significantly different latent Gaussian representations when subjected to an inverse diffusion process using a pre-trained diffusion model. Using these distinctions, they can enhance and accentuate subtle artifacts within generated images, improving the effectiveness of their detection methods. Zhang et al. [57] introduce the Cross Appearance-Edge Learning detector, a tool specifically designed to recognize forgery by capturing both coarse- and fine-grained appearance and edge patterns at multiple levels of granularity. This method effectively merges knowledge from both the appearance and edge domains to thoroughly expose a broad spectrum of forgery artifacts. Nguyen et al.[58] introduce a multi-task learning system with three key branches: classification, heatmap regression, and self-consistency regression. The heatmap regression and self-consistency regression branches are specially designed to pinpoint vulnerable areas within an image, areas that are particularly indicative of blending artifacts, thereby improving the framework’s ability to detect forgeries.
Consistencies-based Methods The field expanded into temporal artifact detection as researchers recognize the importance of dynamic inconsistencies in videos. Li et al.[59] utilize Long-term Recurrent Convolutional Networks to detect inconsistencies around blinking in videos, while Qi et al.[60] introduce DeepRhythm, which analyzes heartbeat rhythms to spot video forgeries. In [61], Zheng et al. propose to leverage temporal coherence for video face forgery detection. Their framework includes two stages: a fully temporal convolution network that enhances temporal feature extraction by reducing the spatial convolution kernel size, and a Temporal Transformer network that explores long-term temporal coherence. Haliassos et al.[62] design a spatio-temporal network to detect semantic inconsistencies in mouth movements within forgery videos.
To further address temporal inconsistencies, various methods have been developed and applied to domains beyond pixel analysis. Choi et al. [63] propose detecting fake videos by analyzing the inconsistent temporal behavior of style latent vectors between real and fake samples. They observe that generated facial videos exhibit distinct temporal changes in these vectors during stable video generation. In their method, Choi et al. employ a StyleGRU module, trained with contrastive learning, to capture the dynamic properties of style latent vectors. The study in [64] introduces latent pattern detection, which identifies temporal inconsistencies across video frames. Haliassos et al.[65] develop a method that employs self-supervised learning to capture detailed video representations by using audio-visual correlations in genuine videos. The detector is then trained on both real and synthetic data to detect facial forgeries and to predict the representations from the initial stage for authentic videos.
Expanding on the identification of spatio-temporal inconsistencies, Pang et al.[66] propose MRE-Net consisted of two main components: Bipartite Group Sampling (BGS) and Multiple Rate Branch (MRB). BGS divides the video into bipartite groups to capture various face motion dynamics, while MRB captures short-term and long-term spatial-temporal inconsistencies, respectively. Wang et al.[67] capture spatio-temporal artifacts in one model based on a disentangled learning strategy, dividing the convolutional weights of a 3D ResNet into spatial-based and temporal-based groups and updating the two groups alternatively during training. Based on the observation that fake videos struggle to maintain natural spatio-temporal consistency under various forgery methods, Zhang et al. [68] introduce the Natural Consistency method for detecting deepfake videos. Their approach leverages CNNs to extract spatial features and Transformers to learn long-range spatiotemporal representations. Additionally, they introduce a Spatial Predictive Module to predict random masked spatial features and a Temporal Contrastive Module to regularize latent distances by shuffling the natural order, both of which effectively enhance detection accuracy.
Several methods focus on using other inconsistencies, such as semantic, identity, and noise patterns, to enhance detection capabilities. Xu et al.[69] introduce the Thumbnail Layout (TALL) strategy to enhance the ability to capture inconsistencies between frames in fake videos. Their following work TALL++[70] is built on top of TALL[69] by designing a graph reasoning block and introducing a consistency loss to better capture semantic-level inconsistency to improve generalizability. TI2Net [71] uses identity inconsistencies in video frames to authenticate clips. It encodes identity information from each frame into vectors and calculates the temporal identity embedding by analyzing the differences between these vectors. The temporal identity embedding effectively captures identity inconsistencies, which is crucial to determining the authenticity of the video input. Based on the assumption that background squares remain unmodified in fake samples, Wang et al. [72] propose utilizing the consistency of noise patterns between the face regions and background patches to differentiate real and fake frames. In [73], Chen et al. focus on detecting compressed deepfake videos, which often present overlapping compression and tampering artifacts.
3.1.1.2   Generalized Detection for GAN/VAE
In the realm of deepfake detection, the generalizability of detection plays an important role in enhancing the adaptability of the detection methods against a wide range of manipulations. In this subsection, we delve into various categories of approaches, such as input level, model level, and learning level methods. Each category provides its own perspectives and techniques to tackle the challenge of generalized deepfake detection.
Input level: Augmentation-based Methods Augmentation-based methods enhance deepfake detection by enriching training datasets with complex pseudo-fake examples. Wang et al. [74] develop an attention-based method that targets key facial regions for data enhancement by adaptively masking areas most likely to be manipulated. Zhao et al. [75] introduce the Inconsistency Image Generator (I2G), creating well-annotated datasets by blending genuine images to avoid common artifacts. Chen et al. [76] further diversify forgery characteristics in training, using adversarial methods to expose the model to a wider array of forgeries and enhance detection capabilities. In further exploration, Shiohara et al. [77] employ “self-blended images” (SBIs) to simulate typical forgery indicators like blending seams. Larue et al. [78] introduce soft discrepancies into synthetic face images, aiding the model in recognizing subtle inconsistencies. Li et al. [190] improve low-quality forgery detection with a “bleach generator” that corrects and enhances input forgeries for better accuracy.
Traditional methods typically involve pixel-level augmentations. However, recent approaches [79, 80] explore augmentations in latent and frequency spaces. Yan et al. [79] use Latent Space Data Augmentation (LSDA) to generate diverse forgery variations, minimizing overfitting and enhancing generalization. FreqBlender [80] focuses on frequency knowledge, employing a Frequency Parsing Network to segment and combine critical frequency components for more authentic pseudo-fakes.
Input Level: Frequency-space based Methods Recently, researchers have expanded beyond RGB space analysis, developing sophisticated techniques that utilize frequency analysis to enhance deepfake detection. Qian et al.[82] introduce the Frequency in Face Forgery Network (F3-Net), utilizing frequency-aware components and statistics to identify forgeries. Chen et al.[83] analyze local region connections using the Multi-scale Patch Similarity Module (MPSM) and RGB-Frequency Attention Module (RFAM) to detect complex patterns and integrate multi-modal data. Li et al.[84] develop an adaptive frequency feature generation module, incorporating a single-center loss (SCL) to distinguish class features effectively.
Building upon these foundational works, subsequent research has introduced advanced approaches with additional complexities. Luo et al.[85] enhance forgery detection with a focus on high-frequency knowledge and cross-modality attention, optimizing data harmonization. Liu et al.[86] present Spatial-Phase Shallow Learning (SPSL) that combines spatial and phase spectrum analysis to spotlight up-sampling artifacts in forged content. Further, Woo et al.[87] present the Attention-based Deepfake detection Distiller (ADD), which applies frequency domain learning and optimal transport theory in knowledge distillation. ADD uses frequency attention distillation to recover high-frequency components lost during compression and multi-view attention distillation to generate multiple attention vectors. Jeong et al.[88] create frequency-level perturbation maps that blur distinctions between genuine and fake images, refining both the detector and perturbation generator. Wang et al.[89] propose the Spatial-Frequency Dynamic Graph method using dynamic graph learning to analyze relation-aware features. Tan et al.[90] develop FreqNet, integrating frequency domain learning with CNN classifiers to enhance detection across various scenarios. Doloriel et al.[91] demonstrate how frequency masking can significantly improve the generalization of deepfake detection, achieving notable performance enhancements.
Model Level: Transformer-Based Approaches Recently, transformers [8] have been employed in both static deepfake detection and temporal deepfake segmentation. Wang et al.[92] provide an overview (as of February 28, 2024) of fourteen works adopting the Transformer architecture for deepfake detection. Building on these advancements, Saha et al. [93] utilize a Transformer-based architecture, integrating a Vision Transformer (ViT) with a Timeseries Transformer (TsT) to focus on manipulated frames, enhancing detection capabilities by analyzing spatial and temporal features. Expanding on the use of transformers, Bai et al. [94] introduce the Action Units Relation Transformer within their Action-Units Relation Learning framework to analyze relationships between facial action units, enhancing forgery detection through AU-specific and agnostic branches.
Further advancing in this direction, Zhang et al.’s UMMAFormer [95] employs the Temporal Feature Abnormal Attention module and a Parallel Cross-Attention Feature Pyramid Network for refined forgery segment prediction. Pellicer et al. [96] propose the Prototype-based Unified Framework, using features from a pre-trained DINOV2 model [191] for prototype learning, aiming to improve deepfake detection through a Prototype Learning layer that clusters and compares prototypes.
Using advances in parameter-efficient tuning methods such as Adaptors[97, 99], Mixture of Experts[100, 99], and LoRA[100, 99], innovative strategies are being proposed to integrate these techniques into large pre-trained Vision-Transformers that serve as foundation models. These strategies elegantly address the data-demanding issue while adapting foundation ViTs to deepfake detection and circumvent distortion of originally learned rich knowledge during pre-training. These techniques efficiently address data demands and prevent distortion of pre-trained knowledge [101]. The DeepFake-Adapter [97] introduces dual-level adapters to ViTs, enhancing anomaly detection without core architecture modifications. Luo et al. [98] develop a Forgery-Aware Adaptive Vision Transformer, employing a dual-branch architecture to boost feature generalization and prevent overfitting. Liu et al.[100] utilize a mixture of low-rank experts to maximize the potential of CLIP-ViT for deepfake detection, aiming to preserve pre-trained knowledge while enhancing the model’s adaptability. Similarly, Kong et al. introduce MoE-FFD [99], a method that incorporates LoRA and Adapter with the Vision Transformer to improve the detection of face forgeries.
Model level: Vision-Language Models Researchers are harnessing the capabilities of pre-trained Vision-Language Models (VLMs) [9], which are proving instrumental not only in universal deepfake detection, but also in enhancing reasoning processes.
Universal Detection: A primary focus has been developing methods that utilize the inherent capabilities of VLMs, such as CLIP [192], to evaluate their effectiveness across various scenarios. To enhance video content analysis, Han et al.[102] propose using the CLIP image encoder, paired with a newly designed side network-based decoder, to effectively extract spatial and temporal cues from video clips. Unlike previous methodologies that directly utilized a frozen CLIP model, they update their CLIP-inspired model parameters by incorporating designed textual labels into the image augmentation process. In [103], Lai et al. propose the Generalized Multi-Scenario Deepfake Detection framework (GM-DF) to enhance the generalization of models trained on multiple datasets. GM-DF uses CLIP to align visual and textual features, hybrid expert modeling for domain-specific extraction, a masked image reconstruction mechanism for detailed forgeries, and a domain-aware meta-learning strategy to improve generalization.
Reasoning: Recent advances have proposed using VLMs to incorporating reasoning to boost interpretability and context understanding in deepfake detection. For example, Tsigos et al. [104] propose a novel framework to evaluate how effectively explanation methods can interpret decisions made by deepfake detectors. This framework focuses on visual data and measures the accuracy of explanation methods by their ability to modify key image regions through adversarial attacks, assessing their impact on the detector’s verdict. The findings underscore the ability of this method to offer meaningful explanations based solely on visual input.
Building on the multimodal capabilities, Sun et al. develop Visual-Linguistic Face Forgery Detection (VLFFD)[105], which employs CLIP to improve face forgery detection. This approach enriches training data with detailed textual prompts and fine-grained images, employing a Coarse-and-Fine Co-training framework that leverages both visual and linguistic knowledge. VLFFD improves forgery identification and provides in-depth interpretative insights into the locations and characteristics of the forgeries, markedly expressing its explanatory power. Zhang et al. [106] reformulate the deepfake problem as a Deepfake Detection Visual Question Answering task, which involves identifying image authenticity and providing explanations. They also create a new dataset and develop a Vision and Language Transformer-based framework that effectively detects deepfakes and generates high-quality explanations.
As Multimodal Large Language Models (MLLMs) have shown significant problem-solving capabilities in various vision fields, recent studies have begun to explore their application in deepfake detection. A recent study by [107] introduces the SHIELD benchmark to evaluate the effectiveness of MLLMs in face forgery detection across both visual and acoustic modalities. Tested under various conditions, including zero-shot and few-shot scenarios, MLLMs demonstrate superior capabilities in interpretability and multimodal reasoning in deepfake detection. In parallel, Shan et al. [108] have validated the efficacy of MLLMs through extensive qualitative and quantitative experiments, highlighting the importance of effective prompt engineering to distinguish between real and AI-generated images effectively. In a recent work[109], Li et al. conduct extensive experiments to evaluate the performance of $13$ well-known language and multi-modal models on the task of detecting fake images and explore how explicit reasoning impacts the accuracy of fake detection with these models.
Learning Level: Advanced Loss Functions While most deepfake detection methods traditionally rely on Softmax loss, Yin et al. [110] challenge the effectiveness of softmax loss for distinguishing subtle distinctions between real and fake images. They propose a method that utilizes both angular and Euclidean spaces to enhance class separability. By decoupling the magnitude and direction of embedding vectors, they introduce margin penalties in both domains to increase the separation between classes, thereby enhancing inter-class separability and intra-class compactness. Zhang et al. [111] further the development of discriminative features using a fractional center loss, designed to cluster real faces closer within the embedding space while distinctly separating them from fake ones. In their subsequent research, Zhang et al. [112] introduce a dual-level center loss approach that incorporates both patch-level and image-level center losses. This innovative method effectively reduces intra-class distances while expanding inter-class distances, significantly enhancing the discriminative capability of feature embeddings at both the patch and image levels.
Learning Level: Disentangled Representation Disentanglement learning addresses the issue that deepfake detectors may learn representations containing both forgery and content information. This mixed representation makes the detector prone to overfitting training data by recognizing content clues such as face identity and background [113] [114]. To mitigate this, disentanglement learning aims to decouple learned characteristics into task-relevant and task-irrelevant groups, ensuring that only task-relevant characteristics are used for subsequent detection.
Liang et al. [113] explore a framework to separate forgery-relevant from content-relevant features, using two dedicated encoders and employing a multi-task strategy and contrastive constraints to enhance the effectiveness of disentanglement. The UCF framework [114] further refines this approach by decoupling forgery-relevant features more distinctly, which aids in better generalization across various forgery types. Lin et al. [115] extend disentanglement learning to include demographic-relevant features, improving both fairness and generalization in deepfake detection. Similarly, Ye et al. [116] utilize an adaptive high-pass module and a two-stage training strategy to enhance the separation of forgery semantics from irrelevant content, aiming to reduce overfitting and improve generalizability. Yang et al. [117] treat different forgery methods as distinct domains and develop the Deep Information Decomposition (DID) framework, which uses attention modules to adaptively separate facial information into forgery-specific categories. This selective focus helps in enhancing detection across different datasets and forgery techniques.
Unlike the aforementioned works that decompose features from a high-level semantic perspective with the aid of multi-task strategy, Ba et al.[118] leverage a disentanglement module to extract multiple non-overlapping local features with the guidance of carefully derived Local Information Loss. These local features are subsequently fused and filtered out task-irrelevant information under the supervision of introduced Global Information Loss.
Learning level: Reconstruction Researchers have innovated methods for identifying deepfakes by harnessing insights gained during the reconstruction of input. Khalid et al.[119] employ a One-class Variational Autoencoder (OC-VAE), trained solely on authentic images, to identify deepfakes by analyzing deviations in reconstruction scores. He et al.[120] utilize image re-synthesis techniques such as super-resolution and denoising to extract residual detection cues. Cao et al.[121] combine reconstruction with metric constraints and classification to effectively distinguish real from forged images. In MARLIN [122], Cai et al. use self-supervised learning to reconstruct accurate facial expressions from extensive online video data. These approaches leverage facial imagery reconstruction as a pivotal element.
Learning level: Manipulation Localization The objective of manipulation localization in deepfake detection is to pinpoint altered pixels in fake samples. This approach not only reveals the specific changes made by attackers but also aids in discerning their motives, which is valuable for informing decisions made by authorities[158].
Li et al.[123] developed Face X-ray, a technique that detects blending boundaries used in face manipulations to differentiate between genuine and altered images. Dang et al.[124] enhance the detection of manipulated face images by using an attention mechanism that refines the processing of feature maps, improving the accuracy of identifying forgeries. Similarly, Shuai et al.[125] employ RGB images and Spatial Rich Mode noise residuals to accurately localize forgery areas by integrating multimodal and multi-scale patch features. Further contributions include Dong et al.[126], who focus on mitigating implicit identity leakage to improve detection generalization with a model that emphasizes both authenticity verification and manipulation localization. Zhang et al.[127] introduce EditGuard, which embeds imperceptible watermarks to enable precise localization of tampered regions in digital content. Recently, Tantaru et al. [129] explore novel pathways in this field. Specifically, Tantaru et al. develop a weakly-supervised technique that improves the localization of manipulated areas in images created through advanced manipulation methods, further refining the ability to detect and understand the nuances of digital forgeries.
Learning level: Identity Discrepancy This approach shifts the focus from simply identifying if a face is real or fake to analyzing whether the facial identity extracted from the input maintains the subject’s biometric traits.
Cozzolino et al. [130] develop a system incorporating a facial feature extractor, a temporal network to identify biometric discrepancies, and a generative adversarial network that predicts subject-specific motions, enhancing the detection of inconsistencies in facial features. Nirkin et al. [131] analyze discrepancies between identity embeddings from the inner(target person) and outer (source person)facial regions. This comparison helps detect forgeries by highlighting inconsistencies in identity representations. Dong et al. [132] introduce the Identity Consistency Transformer, which uses identity vectors to discern differences between the inner and outer facial areas. Their model employs a consistency loss mechanism to align similar identities and differentiate dissimilar ones. Huang et al. [133] leverage implicit identity features of fake faces with two innovative loss functions: the Explicit Identity Contrast loss, which attracts real samples to their true identities and repels fake ones, and the Implicit Identity Exploration loss, a margin-based approach that ensures effective differentiation of genuine and manipulated faces by maintaining appropriate intra-class and inter-class distances.
Learning level: Knowledge Distillation Knowledge distillation techniques [193] have proven effective in enhancing the generalizability of deepfake detection, applicable both at the latent representation and/or prediction levels.
Woo’s Attention-based Deepfake Detection Distiller [134] utilizes both high and low-quality forgery data to identify compressed deepfakes effectively. It employs a frequency attention distiller that aids the student model in recovering lost high-frequency components and a correlation distillation method that enhances pixel correlation preservation using the Sliced Wasserstein Distance metric.
In situations where access to the source domain data is restricted, Kim et al. introduce the Feature Representation Transfer Adaptation Learning (FReTAL) [136]. This method uses knowledge distillation to adapt detectors to new deepfake types efficiently, focusing on latent representations and prediction logits, assuming label availability for new forgeries. Similarly, the CoReD framework addresses sequential domain adaptation challenges by minimizing catastrophic forgetting within a teacher-student model setting [137]. This process ensures the knowledge preservation in both latent representation and prediction levels.
Furthermore, to improve deepfake detection across domains without needing label access for target samples, the study by Lv et al. [138] introduces a bidirectional adaptation strategy. Forward adaptation transfers detection capabilities to new forgeries via supervised training and adversarial feature adaptation, while backward adaptation uses self-distillation to further refine the extracted knowledge.
Learning level: Fine-grained Prediction Recent advancements highlight the importance of granularity and adaptability in facial forgery detection, refining the detection process to keep pace with the evolving complexity and variety of deepfake methods.
In this line of methods, Guo et al. [139] and Guarnera et al. [140] both implement hierarchical approaches to deepfake detection, emphasizing multi-level classification to discern real from synthetic images and further classify the type and source of forgeries. Guo et al.’s method categorizes forgeries from broad groups down to specific techniques, enhancing understanding of the forgeries through a detailed hierarchical analysis. Similarly, Guarnera et al. start with basic real vs. generated classifications, then specify whether the images are from GANs or DMs, and finally identify the precise architecture used. Additionally, the work in [141] on controllable guide-space methods focuses on improving feature discrimination by finely segregating and differentiating forgery domains, thereby boosting generalization capabilities across different forgery scenarios.
3.1.1.3   Detection for Diffusion Model
As diffusion models continue to evolve, their potential misuse for creating counterfeit content has raised significant concerns. The distinctness of their generation process necessitates specialized detectors to discern diffusion-generated images. Given that detection for diffusion models is in its nascent stages, most works are working on images from general domains, of which a subset contains facial forgery images. Nevertheless, these proposed frameworks can be applied to facial forgery images without loss of generality.
The studies in [142] and [143] analyze outputs from DMs and concluded that, despite significant improvements in the quality of the generated results, there are still identifiable clues that can distinguish these from the real samples. Ricker et al. [142] extensively test whether current detection methods designed for GANs generated samples can identify images created by DMs. They find that while these methods initially struggled to tell DM-generated images apart from real ones, retraining them on DM images greatly improved their accuracy, even allowing them to detect images made by GANs. Corvi et al. [143] note that models such as GLIDE [194] and Stable Diffusion [195] produce distinct spectral peaks, which can be effectively used to identify synthetic images. This spectral approach adds another dimension to detection methodologies, harnessing unique signatures inherent in diffusion models.
Lorenz et al. [144] introduce the use of the lightweight multi Local Intrinsic Dimensionality technique, initially developed for detecting adversarial examples, to automatically detect synthetic images and identify the generator networks responsible for their creation. Following a different approach, Cheng et al. [34] present the Edge Graph Regularization method, crafted to enhance the discriminative capabilities of detectors. This technique integrates edge graphs, derived via Sobel operators [196], as a regularization term in the empirical risk formulation. In doing so, the model is encouraged to simultaneously analyze the features of both the original images and the edge graphs, thereby refining its overall analytical precision.
Transitioning from intrinsic analysis to reconstruction accuracy, Wang et al. [145] introduce Diffusion Reconstruction Error (DIRE) that measures the discrepancy between an input image and its reconstructed version using a pre-trained diffusion model. Their method reveals that, while images generated by diffusion models can be accurately reconstructed by the same type of models, real images do not replicate with the same fidelity. Expanding on the concept of reconstruction, Ma et al.[146] introduce a method named Stepwise Error for Diffusion-generated Image Detection (SeDID), addressing the limitations of DIRE[145] by utilizing the intermediate steps in diffusion models. Unlike DIRE that only focuses on the initial time-step information, SeDID captures critical information during the diffusion and reverse diffusion processes, potentially enhancing the accuracy of synthetic image detection. Building on this, Luo et al. [147] capitalize on conducting reconstructions within the latent space, enhancing efficiency in two key ways: it avoids the need for complete image inversion and reconstruction, and it reduces the risk of error accumulation associated with multiple inversion and reconstruction steps.
3.1.1.4   Sequential Deepfake Detection
Based on the observation that the multi-step manipulation of facial components and attributes is becoming possible due to the emergence of facial editing applications, it is crucial to enable the detection of these compound deepfakes beyond one-step manipulation. In response, sequential deepfake detection[148] is proposed, aiming to detect a sequence of facial manipulations. This novel task is significantly more challenging than conventional deepfake detection, requiring a nuanced understanding to detect the facial components and attributes being forged as well as the corresponding manipulation order.
To support this investigation, Shao et al.[148] construct the first sequential deepfake dataset, Seq-DeepFake, which includes sequentially manipulated facial images and corresponding manipulation sequence annotations. They propose SeqFakeFormer, an image-to-sequence model with an image encoder for spatial correlations and a sequence decoder to capture manipulation traces. SeqFakeFormer++[149] further enhances this by incorporating image-sequence reasoning with two modules: Image Sequence Contrastive learning for aligning global image and sequence features, and Image Sequence Matching for fine-grained alignment using a cross-attention mechanism. Hong et al. [150] approach the problem with multi-label ranking for sequence prediction. They decompose the task into binary deepfake classification, manipulation localization, and manipulation order determination. They propose a contrastive multi-instance learning formulation for classification, treating an image as a bag and its patches as instances. For sequential detection, they use multi-label learning with ranking reasoning to reflect the manipulation order by multi-label scores.
3.1.2 Proactive Detection Methods
Proactive detection methods for deepfakes utilize strategies that preemptively counteract deepfake manipulation before it occurs or facilitate the identification of manipulated content afterward. Figure 4 showcases how proactive detection approaches distinguish themselves from passive detection methods. Unlike passive methods, which assess content post-creation, proactive approaches embed specific signals, such as adversarial perturbations, into the original content. These signals aid in verifying the authenticity and integrity if the content is later altered. For effectiveness, these alterations visually resemble the original to evade detection while impairing the functionality of deepfake generation models. In the following subsections, we explore recent proactive detection efforts for GANs and DMs separately.
Figure 4: The comparison between passive and proactive deepfake detection methods. Figure from [154].
3.1.2.1   Proactive Methods for GANs
Researchers have explored using adversarial attack techniques to inject perturbations into the original input, enhancing the resilience of images against unauthorized manipulations. Ruiz et al. [151] develop an approach using versatile adversarial attacks to prevent unauthorized facial modifications, improving the resilience of image translation systems through adversarial training for GANs. Similarly, Yeh et al. [152] use adversarial loss to increase the resistance of images to deepfake manipulations.
To further improve the resilience of perturbations, Huang et al. [153] introduce a poison perturbation generator that improves the effectiveness of perturbations using an alternating training strategy. Wang et al. [154] propose improving the resilience of perturbations by targeting the Lab color space, particularly the decorrelated a and b channels. Expanding defensive strategies, Anejaet al. [155] develop a data-driven method that quickly produces tailored perturbations to redirect manipulation attempts toward a predefined outcome. Huang et al. [156] design the Cross-Model Universal Adversarial Watermark, broadening defense across various images and deepfake models. Asnani et al. [157] and [158] propose using designed real image templates to improve the detection accuracy of manipulation, helping to differentiate between real and fake images.
Recent innovations have suggested conducting source tracing alongside injecting adversarial perturbations into the original input to enhance traceability against manipulations. Innovations by Wang et al. [159] and Wu et al. [160] underscore the importance of tracing original source images to boost detection system credibility. Yu et al. [14] enhance traceability by embedding traceable fingerprints directly into the generator’s parameters, linking outputs uniquely to their creators. Addressing gaps in proactive defenses, Tang et al. [161] introduce the Feature-Output ensemble UNiversal Disruptor, which disrupts in the latent space and employs a gradient-ensemble algorithm to provide a universal attack pattern against multiple deepfake models. In [162], Zhang et al. introduce Dual Defense, which embeds an invisible watermark in target faces to disrupt malicious face swapping while preserving traceability. This approach uses a watermark embedding network that balances invisibility, adversariality, and traceability through perceptual adversarial encoding strategies. In a recent study, Wu et al. [163] demonstrate that traditional watermarking methods, when applied to deepfake detection, can interfere with forgery detection. To address this issue, they propose AdvMark, a plug-and-play solution that retains the ability to extract watermarks for provenance tracking.
3.1.2.2   Proactive Methods for DMs
Recent pioneering strategies in proactive deepfake detection for diffusion models, not specifically designed for facial forgery, include Cui et al.’s DiffusionShield [15], which enhances watermark uniformity, Fernandez et al.’s method [164] for embedding invisible watermarks in Latent Diffusion Models, and Wang et al.’s behavioral modification approach [165] to ensure traceability of protected content. These innovative methods establish foundational practices for applications in facial forgery detection.
Expanding the scope, Zhao et al. [166] examine the effectiveness of watermarking within diffusion models, with Cao et al. [167] extending proactive security measures to audio models. To enhance model transferability, Shim et al. [168] disrupt the latent encoding phase of deepfake generation using Latent Ensemble ATtack, employing a Normalized Gradient Ensemble strategy to optimize various disruptions.
3.2 Multi-modal Deepfake Detection
The concurrence of fake content in multiple modalities leads to more convincing deepfakes compared to those visually manipulated ones. In such context, it becomes imperative to develop the capability to detect these intricate multi-modal deepfakes. In this subsection, we provide a comprehensive investigation of the more intricate and challenging multi-modal scenario, where auditory or textual content can also be falsified along with the presence of visual manipulation.
3.2.1 Audio-Visual Deepfake Detection
Audio-visual deepfakes involve the manipulation of either or both audio and visual streams in a video. The combination of manipulations on these two modalities potentially raises more serious ethical and social concerns, for example, manipulating politicians to make false claims convincingly by delivering a lip-synced video with fake audio.
The heterogeneous nature and the modality gap between audio and video make multi-modal audio-visual deepfake detection a more challenging task. Existing explorations on this task commonly design two-stream framework to incorporate multi-modal inputs. As shown in Figure 5, the efforts can be categorized into three main paradigms: independent learning, joint learning, and matching-based learning.
(a) Independent learning for audio-visual deepfake detection.
(b) Joint learning for audio-visual deepfake detection.
(c) Matching-based learning for audio-visual deepfake detection.
Figure 5: Three main paradigms for multi-modal audio-visual deepfake detection.
Independent Learning The independent learning paradigm is a straightforward approach in audio-visual deepfake detection, where each modality is processed separately. This method then combines the results from each stream for a final decision [10][11][169][170].
As a baseline method presented in [10], two independent streams of convolutional neural network are proposed separately for detecting audio and visual deepfakes, and the final result of multimodal detection is obtained by combining the audio and visual predictions. In a preliminary attempt to evaluate the FakeAVCeleb[5] benchmark, Khalid et al.[11] ensemble predictions from independently learned uni-modal detectors. Furthermore, Hashmi et al.[169] devise an audio-visual network besides uni-modal networks and adopt ensemble learning to make the final decision based on predictions from audio, video, and audio-visual networks. Similarly, AVFakeNet[170] leverages Swin-Transformer [197] as the backbones for both audio and visual branches for feature extraction and then integrates the result from both branches for the multi-modal prediction.
Despite its intuitive nature, the paradigm of independent learning could overlook the intrinsic relationships between audio and visual data, which can be critical in accurately detecting deepfakes.
Joint Learning This approach integrates audio and visual modalities to leverage multi-modal information for enhanced deepfake detection. The general network architecture of the joint learning paradigm encompasses two modal-specific branches taking as input corresponding modality content for feature extraction and a fusion module to integrate features of different modalities for audio-visual detection. The feature fusion module in this paradigm aids to mine the inter-modal relationship. Based on at which stage the modal-specific features are fused together, joint learning-based detection algorithms can be further divided into intermediate and late fusion manners. Moreover, some existing late fusion-based audio-visual detectors leverage multi-task strategy or impose regularization to facilitate the learned multi-modal representations.
Intermediate Fusion: Under the assumption that simply combining the high-level abstract features extracted from both streams is insufficient to mine the inter-modal information, Zhou et al.[10] initially propose the joint audio-visual deepfake detection, handling cases in which either or both visual and auditory modalities have been manipulated. In their method, a sync-stream is designed to fuse the intermediate features from distinct modalities by cross-attention mechanism. Yang et al.[45] employ a Transformer architecture, where audio and visual inputs go through a temporal-spatial encoder, and corresponding features of both modalities interact with each other via a decoder with a bidirectional cross-attention module. Wang et al.[171] incorporate inspiration from how the biological system perceives the multi-modal environment. Marcella et al. [172] extract audio and visual features separately, followed by computing distance and attention maps between the audio and spatial visual features. To address the influence of irrelevant visual regions, such as hair and background, an attention map is applied to diminish their impact on the final classification. In this study, the intermediate audio and visual features attend to each other at all stages and are concatenated at the last few stages for further multi-modal detection.
Late Fusion: This strategy only merges the audio and visual representations in the final stages of feature extraction. Techniques such as concatenation or addition are commonly used to merge modalities in late fusion [11][169][173][174][175][35][13]. Various advanced strategies, including the attention mechanism[176][177] and the MLP mixer layer[12] are applied to enhance effective fusion.
While benchmarking the FakeAVCeleb[5] dataset, Khalid et al. [11] adapt multi-modal models from different tasks to audio-visual deepfake detection, in which the modal-specific features are integrated by concatenation. The audio-visual network presented in [169] uses a visual stream and an audio stream in parallel, taking as input the extracted video frames and mel-spectrogram of audio respectively, and the feature maps from two streams are then concatenated. DF-TransFusion[173] further explores this paradigm by using two branches to extract facial features by using self-attention for facial feature extraction and cross-attention for lip-audio synchronization, with a linear layer integrating the concatenated signals for final prediction.
Further, Liu et al. [175] propose a tri-branch framework to handle audio, video, and multi-scale frame inputs, where the extracted features are weighted based on inter-modal interactions and concatenated. Moreover, this method also designs Forgery Clues Magnification Transformer to improve feature extraction of each branch and Distribution Difference based Inconsistency Computing module to align information of audio and video modalities.
Dynamic weighted addition has proven effective in fusion strategies for multi-modal tasks. While addressing the temporal audio-visual deepfake localization task, Cai et al.[35] calculate the weighted sum of uni-modal boundary maps from corresponding branches as the fused boundary map, where the weights are dependent on uni-modal boundary maps themselves. Similarly, Wang et al.[13] leverage dynamic weighted averaging as late fusion approach in the proposed framework. They fuse modal-specific information via weighted averaging the classification tokens in both uni-modal features, which are extracted by a two-stream transformer. Notably, the weights in this method are calculated by inter-modal attention mechanism.
More sophisticated than simple concatenation or addition, attention mechanisms and MLP mixer layers offer advanced alternatives for multi-modal feature fusion. Sree et al.[177] employ a multi-head self-attention module after concatenating features of different modalities. The attention mechanism further enhances the inter-modal interaction and benefits subsequent multi-modal detection. More than the attention mechanism, Zou et al.[176] feed the concatenated audio-visual features into Transformer layers to fuse information from both modalities. Alternatively, Raza et al.[12] employ an MLP mixer layer that not only enhances intra-modal features but also facilitates the integration of inter-modal information by passing concatenated modal-specific features through an inter-modality MLP mixer layer.
Multi-task: While both intermediate and late fusion-based approaches focus on using fused information solely for multi-modal detection of audio-video pairs, there is a concern that they neglect the fine-grained differences between various types of audio-visual deepfakes. For example, scenarios where the video is real but the audio is fake, or vice versa, can confound feature extraction processes within each modality-specific branch[174]. In response, some works exploit a multi-task learning strategy to address the problem by simultaneously performing uni-modal and multi-modal detection[174, 10, 45, 12, 178].
The networks proposed in [10, 45] are tasked with three binary classification objectives: uni-modal audio deepfake detection, uni-modal visual deepfake detection, and multi-modal detection based on the fused signals. Besides simultaneously introducing uni-modal detection tasks, Muppalla et al.[174] make the multi-modal detection more fine-grained, changing the original binary classification to a four-class one. Similarly, Raza et al.[12] reformulate multi-modal audio-visual deepfake detection as a multi-label classification task. They modify the classification head to incorporate fused signals and simultaneously perform classification tasks for both modalities. The multi-task strategy employed in these studies requires more fine-grained labels of each modality to provide supervision during training.
Unlike those incorporating fine-grained uni-modal detection tasks, Yu et al.[178] design an auxiliary task from a disparate perspective, emphasizing learning content-based cross-modal correlations. In this work, Yu et al. introduce a correlation distillation task, where the student detection network learns to model cross-modal correlation from well-trained audio and visual speech recognition models. This explicitly learned cross-modal correlation based on content helps to prevent overfitting.
Enhanced Representation: This line of work learns more discriminative representations for detection during training. To this end, various strategies including regularization[45][175][176] and disentanglement[177] are employed.
Yang et al.[45] apply a contrastive loss and an additive angular margin loss to make the fused multi-modal features of manipulated and genuine samples more discriminative. Likewise, Zou et al.[176] also impose the contrastive constraint in cross-modal cases, thus the audio and visual features become more similar when they are both pristine. Besides, Zou et al.[176] further introduce within-modal regularization to align uni-modal features , ensuring the uni-modal features are discriminative for corresponding uni-modal detection. Liu et al.[175] also leverage contrastive loss [198] to impose regularization on the relationship between audio and video features based on Jensen-Shannon divergence [199].
Mis-AvoiDD [177] improves multi-modal deepfake detection by decoupling representations into modal-invariant and modal-specific subspaces. It employs a modal-invariant encoder and two modal-specific encoders to extract common and specific features. This method minimizes discrepancies between modal-specific features in the invariant subspace and uses orthogonal loss to maintain distinctness between subspaces, ensuring effective feature alignment.
Matching-based Learning Distinct from fusion-based learning, matching-based learning leverages the natural synchronization found in genuine audio-visual content, which is typically absent in synthetic pairs[179, 180]. This approach focuses on the degree to which the audio and visual streams consistently match, rather than using traditional classification techniques. In this paradigm, networks employ modality-specific feature extractors and are trained on authentic content via contrastive learning, enhancing their generalization capabilities across various manipulation.
Cheng et al.[179] propose Voice-Face matching detection based on the widely verified observation of the homogeneity between one person’s face and voice. The detection network involves only voice and face feature extractors, which are trained by contrastive InfoNCE[200] loss. At inference, the predicted genuineness of the audio-visual pair is determined by the consistency of features from these two modalities. Similar to [179], Tian et al.[180] evaluate the cross-consistency between audio and visual features to compute the matching score, which is compared to a threshold to obtain multi-modal detection results. The method addresses the trade-off observed in deepfake generation between temporal identity consistency and the synchronization of face motion with audio-visual elements by incorporating temporal identity intra-consistency. This method merely required self-supervised learning on real videos by a contrastive learning. Furthermore, Marcella et al.[181] introduce a novel loss that estimates the distance between the 1st-order statistics of audio and visual feature distributions. They maximize this loss for fake data and minimize it for real data, complementing the feature-distance loss.
Others There are several existing works on audio-visual deepfake detection leveraging disparate approaches and frameworks from the aforementioned paradigms.
Feng et al. [182] approach the task from an anomaly detection perspective, requiring training only on real videos in a self-supervised manner. Initially, an audio-visual network is trained to estimate inter-modal synchronization scores, used to extract heuristically designed features. An auto-regressive model then learns the distribution of these features. POI-Forensics [183] focuses on identity consistency rather than cross-modal inconsistency. In this work, Cozzolino et al. argue that manipulation perturbs the portrayed identity’s characteristics. This method requires a reference set of at least 10 videos of the person of interest. The similarity index between the target and reference videos is used to determine the prediction. Haq et al. [184] introduce biological signal emotion psychology to improve deepfake detection. Their neurosymbolic framework performs intra-modal emotion transition analysis and inter-modal reasoning based on extracted emotions from each modality. This framework does not require training on deepfake datasets, relying on psychology and emotion knowledge. Yu et al. [185] propose a two-stage framework for detecting audio-visual deepfakes. The Predictive Visual-audio Alignment Self-supervision stage uses augmented views of visual signals interacting with audio information to obtain VA representations. The designed Multi-modal Deepfake Detection stage concatenates these representations for multi-modal prediction, applying cross-entropy, contrastive, and auxiliary losses. AVFF [186] also employs a two-stage framework. In the first stage, audio-visual representation is learned in a self-supervised manner from real videos using the A2V and V2A networks. These modules are fine-tuned in the second stage, followed by a classifier network. The representation learned in the first stage is highly discriminative for deepfake detection, demonstrating the effectiveness of the proposed audio-visual representation learning paradigm.
3.2.2 Text-Visual Deepfake Detection
Recent advances in text-visual deepfake detection have introduced sophisticated methods to address manipulations on text and visual modalities. Figure 6 illustrates the framework for this multi-modal detection and grounding task, which is conducted in [6].
Figure 6: HAMMER[6] for text-visual deepfake detection and grounding. Figure from [6].
Shao et al. [187] introduce the DGM4 dataset, featuring samples manipulated through distinct image and text techniques. They propose the Hierarchical Multi-modal Manipulation Reasoning Transformer (HAMMER), which enhances cross-modal semantic alignment with a manipulation-aware contrastive loss, demonstrating its effectiveness on this comprehensive dataset.
Liu et al. [188] incorporate frequency domain analysis into RGB image evaluations using discrete wavelet transforms [201] to capture forgery artifacts more effectively. Their approach integrates image and frequency data through a specialized encoder and a unified decoder, streamlining the optimization process. Additionally, the authors utilize the unified decoder featuring symmetric cross-modal interaction modules for gathering modality-specific forgery information and a fusing interaction module to aggregate both visual and textual features.
Wang et al. [189] focus on modality-specific features for detecting and grounding multi-modal manipulations. To simultaneously preserve multi-modal alignment capability and exploit modality-specific features, the framework is devised with dual-branch cross-attention for bi-directional interaction between modalities and decoupled fine-grained classifiers, which perform modality-specific fine-grained manipulation classification task based on features of corresponding modality instead of leveraging an integrated token for both modalities in HAMMER[187]. Furthermore, an implicit manipulation query for each modality is proposed to facilitate reasoning for grounding using learnable queries.
4 Challenges and Future Directions
As discussed in previous sections, deepfake detection has achieved significant progress in recent years. However, various challenges persist due to the complexity of deepfake detection. In this section, we discuss these challenges and present potential avenues for future research.
Performance Generalization Across Scenarios Current deepfake detection methods show limited effectiveness in generalized detection across different datasets, manipulation methods, and post-processing techniques. Variations in these factors can cause significant statistical distribution differences between training and testing sets. Previous research on domain adaptation [202] and domain generalization [203] indicates that ignoring these differences can lead to dramatic performance drops in practical scenarios. Additionally, adversarial attacks challenge the robustness of the models. Considering these challenges, developing specific methods for generalized and robust deepfake detection is a crucial direction for future research.
Alignment Among Multi-modalities The transition from single-modal to multi-modal deepfake detection marks a critical shift in the field. Traditional detection techniques, which primarily focus on single modalities like video or audio, are increasingly inadequate as deepfake generation technologies evolve to incorporate multiple modalities, such as audio-visual or text-visual manipulations. This complexity requires advanced detection strategies that can handle the integration of various modalities. Recent research underscores the potential of exploiting correlations between different modalities to enhance detection accuracy. Future efforts should aim to further develop multi-modal detection frameworks, emphasizing improved synchronization and a deeper understanding of inter-modal dynamics.
Expanding Deepfake Detection with LLMs A recent study [17] has explored using LLMs for generation and detection. We believe that integrating LLMs into deepfake detection can be a promising direction, leveraging their powerful capabilities in detection and explainability. For example, LLMs can analyze subtleties in the text modality that may accompany deepfake videos, such as inconsistencies between spoken words and lip movements. Exploiting these discrepancies can enhance the analysis. Furthermore, LLMs can automatically generate explanations about detected deepfakes, detailing the reasons for the classification and providing evidence, which is crucial for trust and verification purposes.
Privacy-Aware Facial Deepfake Detection Ensuring data privacy in facial deepfake detection is critical due to the sensitive nature of the data and increasing privacy regulations. Most prior works use centralized training, aggregating facial data and increasing the risk of data leakage. Future research should explore privacy-aware methods like federated learning [204]. Federated learning allows training across decentralized devices without data exchange, preserving user privacy [205, 206, 207]. These approaches can enhance compliance with privacy regulations while maintaining high detection performance.
5 Conclusion
This paper provides a thorough review of recent developments in deepfake detection, highlighting research from leading conferences and journals over the past three years. It starts by clearly defining the scope of deepfake detection, followed by an in-depth discussion of datasets and evaluation metrics that address various modalities and generative models. The study methodically reviews and categorizes the principal deepfake detection methods, aiming to help researchers quickly understand current trends. The paper concludes with a discussion on the challenges and prospective future directions in facial deepfake detection.
Acknowledgments Dr. Joey Tianyi Zhou is supported by the National Research Foundation, Prime Minister’s Office, Singapore, and the Ministry of Digital Development and Information, under its Online Trust and Safety (OTS) Research Programme (MCI-OTS-001) and SERC Central Research Fund (Use-inspired Basic Research). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore, or the Ministry of Digital Development and Information.